### v2

# Bangle functions are the generic basis for cluster algebras from punctured surfaces with boundary

Abstract.
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with marked points and non-empty boundary. We prove that for any tagged triangulation $T$ of $\mathbf{\Sigma}$ in the sense of Fomin–Shapiro–Thurston, the coefficient-free bangle functions of Musiker–Schiffler–Williams coincide with the coefficient-free generic Caldero–Chapoton functions arising from the Jacobian algebra of the quiver with potential $(Q(T),W(T))$ associated to $T$ by Cerulli Irelli and the second named author.
When the set of boundary marked points $\mathbb{M}$ has at least two elements, Schröer and the first two authors have shown, relying heavily on results of Mills, Muller and Qin, that the generic coefficient-free Caldero-Chapoton functions form a basis of the coefficient-free (upper) cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$. So, the set of bangle functions proposed by Musiker–Schiffler–Williams over ten years ago is indeed a basis. Previously, this was only known in the unpunctured case.
Key words and phrases:
1. Introduction
1.1. Bases for cluster algebras
Fomin and Zelevinsky introduced around 2001 cluster algebras [17, 18, 19] as a device to study dual canonical bases in quantum groups and the closely related theory of total positivity in complex reductive groups. They expected that in this context all cluster monomials should belong to the dual canonical basis. This question was only recently settled by Kang–Kashiwara–Kim–Oh [30] for the case of generalized symmetric Cartan matrices and by Qin [48] for generalized symmetrizable Cartan matrices. Previously, the first named author together with Leclerc and Schröer [25] had identified the dual of Lusztig’s semicanonical basis with the generic Caldero-Chapoton functions, which also contain all cluster monomials.
Since the set of cluster monomials is always linearly independent [4] this led more generally to the search of bases for (upper) cluster algebras which contain the cluster monomials, and which ideally expand positively with respect to each cluster. Moreover, such ”good” bases are typically parametrized by the by tropical points of the corresponding cluster Poisson variety. By the landmark work of Gross–Hacking–Keel–Kontsevich [28] on theta functions at least one such basis exists under rather weak hypothesis. See the remarkable paper [49] for a thorough discussion.
Parallel to these developments, cluster algebras have found applications in a surprising variety of subjects, see for example the surveys [14], [31],[32], [37].
1.2. Surface cluster algebras and bangle functions.
A prominent class of cluster algebras stems from surfaces with marked points, thanks to works of Fock–Goncharov [13], Fomin–Shapiro–Thurston [15], Fomin–Thurston [16], Gekhtman–Shapiro–Vainshtein [27] and Penner [46].
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a (bordered) surface with marked points in the sense of [15], where $\mathbb{M}$ denotes the set of marked points on the boundary and $\mathbb{P}$ is the set of punctures. See Definition 2.5 for more details. To associate a (coefficient-free) cluster algebra to $\mathbf{\Sigma}$ one considers tagged triangulations of $\mathbf{\Sigma}$. Following [15, Def. 9.6] each tagged triangulation $T$ has an associated skew-symmetric matrix $B(T)$, see also Definition 2.9 for details. The matrix $B(T)$ changes under any flip of an arc in $T$ according to the Fomin–Zelevinsky mutation rule. This allows us already to define a cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{A}(B(T),\mathbf{x})$ which is actually independent of the choice of the triangulation $T$. In fact, a fundamental result of Fomin–Shapiro–Thurston [16, Cor. 6.2] and [15, Thm. 7.11] states that provided $\mathbf{\Sigma}$ is not a closed surface with exactly one puncture, there is a bijection $\gamma\mapsto x_{\gamma}$ between the set of tagged arcs on $\mathbf{\Sigma}$ and the set of cluster variables of $\mathcal{A}(B(T_{0}),\mathbf{x})$, such that the induced assignment $T\mapsto(B(T),(x_{\gamma}))_{\gamma_{\in}T}$ is a bijection between the set of triangulations of $\mathbf{\Sigma}$ and the set of (unlabelled) seeds of $\mathcal{A}(B(T_{0}),\mathbf{x})$, such that whenever two tagged triangulations are related by the flip of a tagged arc, the seeds corresponding to them are related by the mutation of seeds corresponding to the arc being flipped.
Given a surface $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ and a tagged triangulation $T$ of $\mathbf{\Sigma}$, Musiker–Schiffler–Williams associated to each tagged arc (resp. simple closed curve) $\alpha$ on $\mathbf{\Sigma}$, a bipartite graph $G(T,\alpha)$, which they called snake graph (resp. band graph) [43, 44], with edges and tiles naturally labelled by the arcs in $T$ (and the boundary segments of $\mathbf{\Sigma}$). Using these labels, they assigned a weight monomial $w(P):=x(P)y(P)$ to each perfect (resp. good) matching $P$ of $G(T,\alpha)$, thus producing a generating function $\sum_{P}w(P)$ for the perfect matchings (resp. good matchings) of $G(T,\alpha)$. Dividing this generating function by the monomial $\operatorname{cross}(T,\alpha)$ that records the number of crossings of $\alpha$ with each arc in $T$, they define the bangle function
(1.1) $\operatorname{MSW}(G(T,\alpha)):=\frac{\sum_{P}w(P)}{\operatorname{cross}(T,\alpha)}.$
Furthermore, they define the bangle function associated to a lamination $L\in\operatorname{Lam}(\mathbf{\Sigma})$ (see Section 3.3 for details) to be
$\operatorname{MSW}(T,L):=\prod_{\alpha\in L}\operatorname{MSW}(G(T,\alpha))^{m_{\alpha}},$
where $m_{\alpha}$ is the multiplicity of the single laminate $\alpha$ as a member of $L$. The set of bangle functions is
$\mathcal{B}^{\circ}(\mathbf{\Sigma},T):=\{\operatorname{MSW}(T,L)\ |\ \text{$L$ is a lamination of $\mathbf{\Sigma}$}\}.$
Musiker–Schiffler–Williams show in [44] that if the underlying surface $(\Sigma,\mathbb{M},\mathbb{P})$ does not have punctures and has at least two marked points (i.e., $\mathbb{P}=\varnothing$ and $|\mathbb{M}|\geq 2$), then for the coefficient-free cluster algebra $\mathcal{A}(\mathbf{\Sigma})$ associated to $(\Sigma,\mathbb{M},\mathbb{P})$, and for any cluster algebra associated to $(\Sigma,\mathbb{M},\mathbb{P})$ under full-rank extended exchange matrices $\left[\begin{array}[]{c}B(T)\\ B^{\prime}\end{array}\right]$, that (after appropriate specialisations in the sense of Fomin-Zelevinsky’s separation of additions formula [Theorem 3.7, [19]]) both the set of bangle functions and the set of bracelet functions form bases of the associated cluster algebra, each containing the cluster monomials by [43]. They conjecture in [44] that $\mathcal{B}^{\circ}$
forms a basis also in the punctured case.
1.3. Generic Caldero-Chapoton functions.
Let $Q$ be a 2-acyclic quiver with vertex set $Q_{0}=\{1,2,\ldots,n\}$, and $B(Q)=(b_{ij})\in\mathbb{Z}^{Q_{0}\times Q_{0}}$ the corresponding skew-symmetric matrix with with entries
(1.2) $b_{ij}=\lvert\{\alpha\in Q_{1}\mid\alpha:j\rightarrow i\ \text{in}\ Q\}\rvert-\lvert\{\beta\in Q_{0}\mid\beta:i\rightarrow j\ \text{in}\ Q\}\rvert.$
as in [10, Equation (1.4)]. In this context we abbreviate $\mathcal{A}(Q):=\mathcal{A}(B(Q))$ for the corresponding cluster algebra with trivial coefficients.
Recall that for each basic $\mathbb{C}$-algebra $A=\mathbb{C}\langle\!\langle Q\rangle\!\rangle/I$ and a representation $M$ of $A$ we have the $F$-polynomial
$F_{M}(y_{1},\ldots y_{n}):=\sum_{\mathbf{e}\leq\underline{\dim}(M)}\chi(\operatorname{Gr}^{A}_{\mathbf{e}}(M))\mathbf{y}^{\mathbf{e}}\in\mathbb{Z}[y_{1},\ldots y_{n}]$
where $\mathbf{y}^{\mathbf{e}}:=\displaystyle\prod_{i=1}^{n}y_{i}^{e_{i}}$ and $\chi(\operatorname{Gr}^{A}_{\mathbf{e}}(M))$ is the topological Euler characteristic of the quiver Grassmannian of subrepresentations of $M$ with dimension vector $\mathbf{e}$.
Building on this, one may define the Caldero–Chapoton function
$CC_{A}(M):=\mathbf{x}^{\mathbf{g}(M)}F_{M}(\hat{y}_{1},\ldots\hat{y}_{n})\in\mathbb{Z}[x^{\pm 1}_{1},\ldots x^{\pm 1}_{n}]$
where $\hat{y}_{j}:=\displaystyle\prod_{i=1}^{n}x_{i}^{b_{ij}}$ for each $j\in Q_{0}$ and $\mathbf{g}(M)\in\mathbb{Z}^{Q_{0}}$ is the injective g-vector of $M$.
Together with Leclerc and Schröer, the first named author explored in [25] the algebraic geometry of (affine) varieties of representations of Jacobi-finite non-degenerate quivers with potential $(Q,S)$ which appear in algebraic Lie theory. They isolated a class of irreducible components of these representation spaces with good geometric-homological properties and called them strongly reduced irreducible components. Later on, the adjective used was changed to generically $\tau^{-}$-reduced, or simply, $\tau^{-}$-reduced, cf. [23]. Roughly speaking, a component $Z\subseteq\operatorname{rep}(\mathcal{P}(Q,S),\mathbf{d})$ is $\tau^{-}$-reduced if the codimension in $Z$ of a top dimensional $\operatorname{GL}_{\mathbf{d}}(\mathbb{C})$-orbit $\mathcal{O}\subseteq Z$ is equal to the minimum value taken on $Z$ by Derksen-Weyman-Zelevinsky’s (injective) $E$-invariant. See [6, Sec. 3 and 5] for more details.
It is easy to see that each $\tau^{-}$-reduced irreducible component $Z$ has an open dense subset where the constructible function $M\mapsto CC_{\mathcal{P}(Q,S)}(M)$ takes a constant value, which is then denoted $CC_{\mathcal{P}(Q,S)}(Z)$. Resorting to decorated representations and spaces of decorated representations in order to be able to hit initial cluster variables through the Caldero-Chapoton function, the set of generic Caldero-Chapoton functions is
[…]
It is in fact quite easy to see with the help of [10] that for each Jacobi-finite non-degenerate quiver with potential $(Q,S)$ the set $\mathcal{B}_{\mathcal{P}(Q,S)}$ contains all cluster monomials of the cluster algebra $\mathcal{A}(Q)$, and that all elements of $\mathcal{B}_{\mathcal{P}(Q,S)}$ belong to the upper cluster algebra $\mathcal{U}(Q)$ corresponding to $Q$. If the set $\mathcal{B}_{\mathcal{P}(Q,S)}$ is a basis of the (coefficient-free) upper cluster algebra then it is called the generic basis of $\mathcal{U}(Q)$.
One of the main results of [25] is, that in the setting of unipotent cells for symmetric Kac-Moody groups, the generic Caldero–Chapoton functions can be identified with the dual of Lusztig’s semicanonical basis [38], and that this basis contains in particular all cluster monomials. This led the authors of [25] to conjecture that the set $\mathcal{B}_{\mathcal{P}(Q,S)}$ might be a basis of the upper cluster algebra in more situations beyond the algebraic Lie theoretic context.
The set $\mathcal{B}_{\mathcal{P}(Q,S)}$ does not change (up to cluster automorphism) under mutations as a subset of $\mathcal{U}(Q)$. This was shown by Plamondon when $(Q,S)$ is Jacobi-finite non-degenerate [47], and by Schröer and the first two authors when $(Q,S)$ is arbitrary non-degenerate [22].
1.4. Statement of the Main Result
Let $\mathbf{\Sigma}$ be a surface with marked points. To each tagged triangulation $T$ of $\mathbf{\Sigma}$ we associate a quiver $Q^{\prime}(T)$ such that $B(Q^{\prime}(T))=B(T)$. This convention allows us to state and prove our results without the roundabout with dual CC-functions as for example in [23]. Schröer and the first two named authors showed [22, Thm. 1.18], relying heavily on results of Mills [40], Muller [41, 42] and Qin [49], that when $(Q,S)=(Q^{\prime}(T),S^{\prime}(T))$ for some triangulation $T$ of a surface $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ with at least two marked points on the boundary, the set of generic CC-functions $\mathcal{B}_{\mathcal{P}(Q,S)}$ is indeed a basis for the (upper) cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$, where $S^{\prime}(T)$ is the potential defined by Cerulli Irelli and the second author in [5]. Let us write from now on $A(T)$ for the corresponding complex Jacobian algebra $\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T),W^{\prime}(T))$. Note that due to our convention here, the algebra $A(T)$ is opposite to the one which was considered for example in [24]. We are now ready to state our main result, Theorem 6.3, in a slightly simplified version.
Theorem 1.1.
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with marked points such that $|\mathbb{M}|\geq 2$. For each tagged triangulation $T$ of $\mathbf{\Sigma}$, the set $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)$ of coefficient-free bangle functions is equal to the generic basis $\mathcal{B}_{A(T)}$ of the coefficient-free (upper) cluster algebra $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$. In particular, the set $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)$ is a basis of the cluster algebra.
Remark 1.2.
(1) The proof of our theorem relies heavily on [24] and our Combinatorial Key Lemma 5.3. These arguments imply $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)=\mathcal{B}_{A(T)}$. See Section 1.5 for more details. By the above mentioned result from  [22, Thm. 1.18], the set $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)$ is indeed a basis of the upper cluster algebra.
(2) As we have mentioned already, Musiker–Schiffler–Williams showed in [44] that in the special case $\mathbb{P}=\emptyset$ the bangle functions form a basis of $\mathcal{A}(\mathbf{\Sigma})$, and they conjectured in Appendix A of loc. cit. that the restriction to $\mathbb{P}=\emptyset$ can be removed.
(3) Schröer and the first two named authors showed in [23, Thm. 11.9] this theorem for the special case $\mathbb{P}=\emptyset$. Their proof relies heaavily on the fact that for $\mathbb{P}=\emptyset$ the Jacobian algebra $A(T)$ is gentle, which allows for quite direct calculations. In our more general situation the algebras $A(T)$ are still tame [21], but the description of their indecomposable representations is in general unknown. See for example [11] for partial results into this direction. Our approach here exploits the use of mutations in the spirit of [10].
(4) Gross–Hacking–Keel–Kontsevich construct in their ground breaking paper [28] a theta basis for a large class of (upper) cluster algebras with skew-symmetrizable exchange matrix $B$. This basis has remarkable positivity properties.
Recently, Mandel–Qin [39] have shown that when $B=B(T)$ for some triangulation of a surface $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$, Gross–Hacking–Keel–Kontsevich’s theta basis is equal to the set of bracelet functions $\mathcal{B}(\mathbf{\Sigma})$ from [44], thus proving, through a combination with [28], that the set of bracelet functions is indeed a basis. The strategy of their proof is surprisingly similar to ours.
1.5. Structure of the paper
The paper is organized as follows. In Section 2 we provide the basics about mutations and cluster algebras from surfaces. In Section 3 we summarize the main constructions and results from [24] upon which we will rely in Section 6, the most relevant one being [24, Proposition 2.8] (stated as Proposition 3.8 below), which gives a precise recursive formula for the change that the generic projective $g$-vector of any $\tau$-reduced component undergoes under mutation of $\tau$-reduced components.
It should be stressed that, since elements of $\mathcal{B}^{\circ}$ were defined directly from the associated snake or band graphs, then, a priori, this viewpoint says nothing about whether $\mathcal{B}^{\circ}$ is canonical in a cluster-theoetrical sense. That is, whether or not $\mathcal{B}^{\circ}$ is independent of the choice of triangulation, up to a cluster isomorphism.
Throughout this paper, our approach works by flipping the perspective of [43] — we fix a curve on the surface and alter the underlying triangulations through the process of cluster mutation. Since we wish to work with principal coefficients, extra care is needed here. Indeed, we consequently require a framework that describes how cluster variables transform from one triangulation to the next, with respect to principal coefficients at those triangulations. This is achieved through the following lemma:
Lemma 1.3 (Combinatorial Key Lemma).
Let $T=\{\tau_{1},\ldots,\tau_{n}\}$ be a tagged triangulation of $\mathbf{\Sigma}$, $\tau_{k}$ a tagged arc belonging to $T$, and $\alpha$ a simple closed curve. We set $T^{\prime}$ to be the tagged triangulation obtained from $T$ by flipping $\tau_{k}$. The following equality holds:
(1.3) $(y_{k}+1)^{h_{k}}F_{G(T,\alpha)}(y_{1},\ldots,y_{n})=(y_{k}^{\prime}+1)^{h_{k}^{\prime}}F_{G(T^{\prime},\alpha)}(y_{1}^{\prime},\ldots,y_{n}^{\prime})$
where
  • $(\mathbf{y}^{\prime},B^{\prime})\in\mathbb{Q}(y_{1},\ldots,y_{n})$ is obtained from $(\mathbf{y},B(T))$ by mutation at $k$,
  • the snake $\mathbf{h}$-vector $\mathbf{h}_{G(T,\alpha)}$ is defined by setting
    $u^{h_{i}}:={F_{G(T,\alpha)}}_{{|}_{\operatorname{Trop}(u)}}(u^{[-b_{i1}]_{+}},\ldots,u^{-1},\ldots,u^{[-b_{in}]_{+}}),$
    where $u^{-1}$ is in the $i^{th}$ component. The snake $\mathbf{h}$-vector $\mathbf{h}_{G(T^{\prime},\alpha)}$ is defined analogously with respect to $T^{\prime}$ and $\mu_{k}(B(T))=B(T^{\prime})$.
Moreover, the snake $\mathbf{g}$-vector $\mathbf{g}_{G(T,\alpha)}=(g_{1},\ldots,g_{n})$ satisfies
(1.4) $g_{k}=h_{k}-h_{k}^{\prime}$
and $\mathbf{g}_{G(T,\alpha)}=(g_{1},\ldots,g_{n})$ is related to $\mathbf{g}_{G(T^{\prime},\alpha)}=(g_{1}^{\prime},\ldots,g_{n}^{\prime})$ by the following rule:
(1.5) $g_{j}^{\prime}:=\left\{\begin{array}[]{ll}-g_{k},&\text{if $j=k$};\\ \\ g_{j}+[b_{jk}]_{+}g_{k}-b_{jk}h_{k},&\text{if $j\neq k$}.\\ \end{array}\right.$
As an immediate corollary, we obtain that the set of bangle functions $\mathcal{B}^{\circ}(\mathbf{\Sigma})=\mathcal{B}^{\circ}(\mathbf{\Sigma},T)$ is, up to cluster automorphism, independent of the choice of $T$. Note that a version of this lemma first appeared in the work of Derksen-Weyman-Zelevinsky [[10], Lemma 5.2] which was restricted the setting of cluster variables — our result shows this lemma holds on all elements of the basis $\mathcal{B}^{\circ}$.
In Section 4 we show that for simple closed curve $\alpha$, there exists a tagged $T_{\alpha}$ for which one can explicitly compute a (decorated) representation $\mathcal{M}(T_{\alpha},\alpha)$ such that
  • the projective $g$-vector $\mathbf{g}_{A(T_{\alpha})}(\mathcal{M}(T_{\alpha},\alpha))$ is equal to the vector of dual shear coordinates $\operatorname{Sh}_{T_{\alpha}}(\alpha)$ and to the snake $g$-vector $\mathbf{g}_{G(T_{\alpha},\alpha)}$ coming from the snake or band graph $G(T_{\alpha},\alpha)$;
  • the representation-theoretic (dual) $F$-polynomial $F_{M(T_{\alpha},\alpha)}$ is equal to the snake $F$-polynomial $F_{G(T_{\alpha},\alpha)}$ coming from the snake or band graph $G(T_{\alpha},\alpha)$;
  • the orbit closure of $\mathcal{M}(T_{\alpha},\alpha)$ is a $\tau$-reduced component $Z_{T_{\alpha},\alpha}$ and any decorated representation mutation-equivalent to $\mathcal{M}(T_{\alpha},\alpha)$ is generic, for the values of the projective $g$-vector, the representation-theoretic (dual) $F$-polynomial and Derksen-Weyman-Zelevinsky’s $E$-invariant, in the corresponding $\tau$-reduced component.
In Section 5 we show that every time a flip of tagged triangulations is applied, the changes undergone by the snake $g$-vectors and snake $F$-polynomials of the band graphs constructed from $\alpha$, obey the same recursive formulas that govern the changes undergone by the projective $g$-vectors and representation-theoretic (dual) $F$-polynomials. In Section 6 we combine the results from Sections 4 and 5 to deduce our main result. In Section 7 we make some remarks about the extent to which the representations that yield the bangle functions have been really explicitly computed, and state a couple of open problems.
2. Preliminaries
2.1. Cluster algebras
This section provides a brief review of (skew-symmetric) cluster algebras of geometric type. Let $n\leq m$ be positive integers. Furthermore, let $\mathcal{F}$ be the field of rational functions in $m$ independent variables. Fix a collection $X_{1},\ldots,X_{n},x_{n+1},\ldots,x_{m}$ of algebraically independent variables in $\mathcal{F}$. We define the coefficient ring to be $\mathbb{ZP}:=\mathbb{Z}[x^{\pm 1}_{n+1}\ldots x_{m}^{\pm 1}]$.
Definition 2.1.
A (labelled) seed consists of a pair, $(\mathbf{x},\mathbf{y},B)$, where
  • $\mathbf{x}=(x_{1},\ldots x_{n})$ is a collection of variables in $\mathcal{F}$ which are algebraically independent over $\mathbb{ZP}$,
  • $\mathbf{y}=(y_{1},\ldots y_{n})$ where $y_{k}=\displaystyle\prod_{j=n+1}^{m}x_{j}^{b_{jk}}$ for some $b_{jk}\in\mathbb{Z}$,
  • $B=(b_{jk})_{j,k\in\{1,\ldots,n\}}$ is an $n\times n$ skew-symmetric integer matrix.
The variables in any seed are called cluster variables. The variables $x_{n+1},\ldots,x_{m}$ are called frozen variables. We refer to $\mathbf{y}$ as the choice of coefficients.
Definition 2.2.
Let $(\mathbf{x},\mathbf{y},B)$ be a seed and let $i\in\{1,\ldots,n\}$.
We define a new seed $\mu_{i}(\mathbf{x},\mathbf{y},B):=(\mathbf{x}^{\prime},\mathbf{y}^{\prime},B^{\prime})$, called the mutation of $(\mathbf{x},\mathbf{y},B)$ at $i$ where:
  • $\mathbf{x}^{\prime}=(x^{\prime}_{1},\ldots x^{\prime}_{n})$ is defined by
    $x^{\prime}_{i}=\frac{\displaystyle\prod_{b_{ki}>0}x_{k}^{b_{ki}}+\prod_{b_{ki}<0}x_{k}^{-b_{ki}}}{x_{i}}$
    and setting $x_{j}^{\prime}=x_{j}$ when $j\neq i$;
  • $\mathbf{y}^{\prime}$ and $B^{\prime}=(b^{\prime}_{jk})$ are defined by the following rule:
    $b^{\prime}_{jk}=\begin{cases}-b_{jk},&\text{if $j=i$ or $k=i$,}\\ b_{jk}+\max(0,-b_{ji})b_{ik}+\max(0,b_{ik})b_{ji},&\text{otherwise.}\\ \end{cases}$
A quiver is a finite directed (multi) graph $Q=(Q_{0},Q_{1})$ where $Q_{0}$ is the set of vertices and $Q_{1}$ is the set of directed edges. It will often be convenient to encode (extended) skew-symmetric matrices as quivers. We describe this simple relationship in the definition below, which follows the conventions set out in the work of Derksen, Weyman and Zelevinsky [10].
Definition 2.3.
Given a skew-symmetric $n\times n$ matrix $B$ we define a quiver $Q(B)=(Q_{0},Q_{1})$ by setting $Q_{0}=\{1,\ldots,n\}$ and demanding that for any $i,j\in Q_{0}$ there are $[b_{ij}]_{+}$ arrows from $j$ to $i$ in $Q_{1}$.
We generalise the definition to extended $m\times n$ skew-symmetric matrices $\tilde{B}$ in the obvious way.
Definition 2.4.
Fix an $(\mathbf{x},\mathbf{y},B)$. If we label the initial cluster variables of $\mathbf{x}$ from $1,\ldots,n$ then we may consider the labelled n-regular tree $\mathbb{T}_{n}$. Each vertex in $\mathbb{T}_{n}$ has $n$ incident edges labelled $1,\ldots,n$. Vertices of $\mathbb{T}_{n}$ represent seeds and the edges correspond to mutation. In particular, the label of the edge indicates which direction the seed is being mutated in.
Let $\mathcal{X}$ be the set of all cluster variables appearing in the seeds of $\mathbb{T}_{n}$. The cluster algebra of the seed $(\mathbf{x},\mathbf{y},B)$ is defined as $\mathcal{A}(\mathbf{x},\mathbf{y},B):=\mathbb{ZP}[\mathcal{X}]$.
We say $\mathcal{A}(\mathbf{x},\mathbf{y},B)$ is the cluster algebra with principal coefficients if $m=2n$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ satisfies $y_{k}=x_{n+k}$ for any $k\in\{1,\ldots,n\}$.
2.2. Cluster algebras from surfaces
In this subsection we recall the work of Fomin, Shapiro and Thurston [15], which establishes a cluster structure for triangulated orientable surfaces.
Definition 2.5.
A surface with marked points is a triple $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ consisting of a compact, connected, oriented two-dimensional real differentiable manifold $\Sigma$ with (possibly empty) boundary $\partial\Sigma$, a finite set $\mathbb{M}\subseteq\partial\Sigma$ containing at least one point from each boundary component of $\Sigma$, and a finite set $\mathbb{P}\subseteq\Sigma\setminus\partial\Sigma$. It is required that $\mathbb{M}\cup\mathbb{P}\neq\varnothing$. The elements of $\mathbb{M}\cup\mathbb{P}$ are called marked points and the elements of $\mathbb{P}$ are called punctures. If $\mathbb{P}=\varnothing$, it is said that $\mathbf{\Sigma}$ is unpunctured, whereas if $\mathbb{P}\neq\varnothing$, one says that $\mathbf{\Sigma}$ is punctured. For technical reasons we exclude the cases where $\mathbf{\Sigma}$ is an unpunctured or once-punctured monogon; a digon; a triangle; or a once, twice or thrice punctured sphere. In [15, Def. 2.1] the term bordered surfaces with marked points was used instead of the more recent terminology surface with marked points or even marked surface.
Definition 2.6.
An arc of $\mathbf{\Sigma}$ is a simple curve in $\Sigma$ connecting two marked points of $\mathbb{M}\cup\mathbb{P}$, which is not homotopic to a boundary segment or a marked point. We consider arcs up to homotopy (relative to its endpoints) and up to inversion of the orientation.
A tagged arc $\gamma$ is an arc whose endpoints have been ‘tagged’ in one of two ways; plain or notched. Moreover, this tagging must satisfy the following conditions: if the endpoints of $\gamma$ share a common marked point, they must receive the same tagging; and an endpoint of $\gamma$ lying on the boundary $\partial S$ must always receive a plain tagging. In this paper we shall always consider tagged arcs up to the equivalence induced from the equivalence relation for plain arcs.
Definition 2.7.
Let $\alpha$ and $\beta$ be two tagged arcs of $\mathbf{\Sigma}$. We say $\alpha$ and $\beta$ are compatible if and only if the following conditions are satisfied:
  • There exist homotopic representatives of $\alpha$ and $\beta$ that don’t intersect in the interior of $\Sigma$.
  • Suppose the untagged versions of $\alpha$ and $\beta$ do not coincide. If $\alpha$ and $\beta$ share an endpoint $p$ then the ends of $\alpha$ and $\beta$ at $p$ must be tagged in the same way.
  • Suppose the untagged versions of $\alpha$ and $\beta$ do coincide. Then precisely one end of $\alpha$ must be tagged in the same way as the corresponding end of $\beta$.
A tagged triangulation of $\mathbf{\Sigma}$ is a maximal collection of pairwise compatible tagged arcs of $\mathbf{\Sigma}$. Moreover, this collection is forbidden to contain any tagged arc that enclose a once-punctured monogon.
An ideal triangulation of $\mathbf{\Sigma}$ is a maximal collection of pairwise compatible plain arcs. Note that ideal triangulations decompose $\mathbf{\Sigma}$ into triangles, however, the sides of these triangles may not be distinct; two sides of the same triangle may be glued together, resulting in a self-folded triangle.
Remark 2.8.
To each tagged triangulation $T$ we may uniquely assign an ideal triangulation $T^{\circ}$ as follows:
  • If $p$ is a puncture with more than one incident notch, then replace all these notches with plain taggings.
  • If $p$ is a puncture with precisely one incident notch, and this notch belongs to $\beta\in T$, then replace $\beta$ with the unique arc $\gamma$ of $\mathbf{\Sigma}$ which encloses $\beta$ and $p$ in a monogon.
Conversely, to each ideal triangulation $T$ we may uniquely assign a tagged triangulation $\iota(T)$ by reversing the second procedure described above.
Definition 2.9.
Let $T$ be a tagged triangulation, and consider its associated ideal triangulation $T^{\circ}$. We may label the arcs of $T^{\circ}$ from $1,\ldots,n$ (note this also induces a natural labelling of the arcs in $T$). We define a function, $\pi_{T}:\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}$, on this labelling as follows:
$\pi_{T}(i)=\begin{cases}j&\text{if $i$ is the folded side of a self-folded triangle in $T^{\circ}$,}\\ &\text{and $j$ is the remaining side;}\\ i&\text{otherwise.}\\ \end{cases}$
For each non-self-folded triangle $\Delta$ in $T^{\circ}$, as an intermediary step, define the matrix $B_{T}^{\Delta}=(b^{\Delta}_{jk})$ by setting
$b^{\Delta}_{jk}=\begin{cases}1&\text{if $\Delta$ has sides $\pi_{T}(j)$ and $\pi_{T}(k)$, and $\pi_{T}(k)$ follows $\pi_{T}(j)$}\\ &\text{in the clockwise sense;}\\ -1&\text{if $\Delta$ has sides $\pi_{T}(j)$ and $\pi_{T}(k)$, and $\pi_{T}(k)$precedes $\pi_{T}(j)$}\\ &\text{in the clockwise sense;}\\ 0&\text{otherwise.}\\ \end{cases}$
The adjacency matrix $B(T)=(b_{ij})$ of $T$ is then defined to be the following summation, taken over all non-self-folded triangles $\Delta$ in $T^{\circ}$:
$B(T):=\sum\limits_{\Delta}B_{T}^{\Delta}$
Definition 2.10.
Let $T$ be a tagged triangulation of a bordered surface $\mathbf{\Sigma}$. Consider the initial seed $(\mathbf{x},\mathbf{y},B_{T})$, where: $\mathbf{x}$ contains a cluster variable for each arc in $T$; $B(T)$ is the matrix defined in Definition 2.9; and $\mathbf{y}$ is any choice of coefficients. We call $\mathcal{A}(\mathbf{x},\mathbf{y},B(T))$ a surface cluster algebra.
Proposition 2.11 (Theorem 7.9, [15]).
Let $T$ be a tagged triangulation of a bordered surface $\mathbf{\Sigma}$. Then for any $\gamma\in T$ there exists a unique tagged arc $\gamma^{\prime}$ on $\mathbf{\Sigma}$ such that $f_{\gamma}(T):=(T\setminus\{\gamma\})\cup\{\gamma^{\prime}\}$ is a tagged triangulation. We call $f_{\gamma}(T)$ the flip of $T$ with respect to $\gamma$.
Theorem 2.12 (Theorem 6.1, [15]).
Let $\mathbf{\Sigma}$ be a surface with marked points. If $\mathbf{\Sigma}$ is not a once punctured closed surface, then in the cluster algebra $\mathcal{A}(\mathbf{x},\mathbf{y},B(T))$, the following correspondence holds:
$\hskip 65.44133pt\mathcal{A}(\mathbf{x},\mathbf{y},B(T))$ $\mathbf{\Sigma}\hskip 91.04881pt$
[…]
When $\mathbf{\Sigma}$ is a once-punctured closed surface then cluster variables are in bijection with all plain arcs or all notched arcs depending on whether $T$ consists solely of plain arcs or notched arcs, respectively.
3. Laminations as $\tau$-reduced components
3.1. Tame partial KRS-monoids
The simple-minded notion of partial Krull-Remak-Schmidt monoid was introduced in [24] as a convenient abstract framework to state the naturalness of the bijection between laminations and $\tau$-reduced components proved therein.
Definition 3.1.
[24, Definition 2.1] A partial monoid is a triple $(X,e,\oplus)$ consisting of a set $X$, a symmetric function $e:X\times X\rightarrow\mathbb{N}$ and a partially-defined sum :$\oplus:\{(x,y)\in X\times X\mid e(x,y)=0\}\rightarrow X$ such that:
  • (s)
    if $e(x,y)=0$ we have $x\oplus y=y\oplus x$;
  • (00)
    there exists a unique element $0\in X$ with $e(0,x)=0$ and $0\oplus x=x$ for all $x\in X$;
  • (d)
    if $e(y,z)=0$ we have $e(x,y\oplus z)=e(x,y)+e(x,z)$ for all $x\in Z$;
  • (a)
    $(x\oplus y)\oplus z=z\oplus(y\oplus z)$ whenever one side of the equation is defined.
A morphism of partial monoids from $X=(X,e,\oplus)$ to $X^{\prime}=(X^{\prime},e^{\prime},\oplus^{\prime})$ is a function $f:X\rightarrow Y$, such that $e^{\prime}(f(x),f(y))=e(x,y)$ for all $x,y\in X$ and $f(x\oplus y)=f(x)\oplus^{\prime}f(y)$ whenever $e(x,y)=0$.
Remark 3.2.
(1) Suppose $(x_{1}\oplus x_{2})\oplus x_{3}$ is defined. By $(d)$, we have $e(x_{i},x_{j})=0$ for $1\leq i<j\leq 3$, hence $x_{1}\oplus(x_{2}\oplus x_{3})$ is defined.
(2) Suppose that we have $x_{1},x_{2},\ldots,x_{n}\in X$ with $e(x_{i},x_{j})=0$ for all $i<j$, then $x_{1}\oplus x_{2}\oplus\cdots\oplus x_{n}\in X$ is well-defined, and for each permutation $\sigma\in\mathfrak{S}_{n}$ we have
$x_{1}\oplus x_{2}\oplus\cdots\oplus x_{n}=x_{\sigma(1)}\oplus x_{\sigma(2)}\oplus\cdots\oplus x_{\sigma(n)}.$
Definition 3.3.
[24, Definition 2.3] Suppose $X=(X,e,\oplus)$ is a partial monoid.
  • The elements of the set
    $X_{\mathrm{ind}}:=\{x\in X\setminus\{0\}\mid x=y\oplus z\text{ implies }y=0\text{ or }z=0\}$
    are called indecomposable elements of $X$.
  • We say that $X$ is a partial KRS-monoid if every $x\in X$ is equal to a finite direct sum of indecomposable elements, and whenever
    $x_{1}\oplus x_{2}\oplus\cdots\oplus x_{m}=y_{1}\oplus y_{2}\oplus\cdots\oplus y_{n}$
    with $x_{1},\ldots,x_{m},y_{1},\ldots,y_{m}\in X_{\mathrm{ind}}$, necessarily $m=n$ and, moreover, there exists a permutation $\sigma\in\mathfrak{S}_{n}$ such that $y_{i}=x_{\sigma(i)}$ for all $i=1,2,\ldots,n$.
  • We say that $X$ is tame if $e(x,x)=0$ for all $x\in X$.
  • A framing for $X$ is a map $\mathbf{g}:X\rightarrow\mathbb{Z}^{n}$ (for some non-negative integer $n$) such that $\mathbf{g}(x\oplus y)=\mathbf{g}(x)+\mathbf{g}(y)$ for all $x,y\in C$ with $e(x,y)=0$. A framing is faithful if it is an injective function.
  • A framed partial monoid $X=(X,e,\oplus,\mathbf{g})$ is free of rank $n$ if the framing $\mathbf{g}:X\rightarrow\mathbb{Z}^{n}$ is bijective.
Example 3.4.
[24, Example 2.4] Suppose $C$ is a set equipped and $e:C\times C\rightarrow\mathbb{N}$ is a function such that $e(c,c)=0$ for all $c\in C$. Then
$\operatorname{KRS}(C,e):=\{f:C\rightarrow\mathbb{N}\mid c_{1},c_{2}\in\operatorname{supp}(f)\Rightarrow e(c_{1},c_{2})=0\text{ and }\lvert\operatorname{supp}(f)\rvert<\infty\}$
is a tame partial KRS-monoid with
$e(f,g):=\sum_{c,d\in C}f(c)\cdot g(d)\cdot e(c,d),\qquad\text{and}\qquad(f\oplus g)(c):=f(c)+g(c).$
Notice that
$\operatorname{KRS}(C,e)_{\mathrm{ind}}=\{\delta_{c}:C\rightarrow\mathbb{N}\mid c\in C\},\quad\text{where}\quad\delta_{c}(d):=\delta_{c,d}\quad\text{(Kronecker's \emph{delta})}.$
With this notation we have $f=\oplus_{c\in C}\delta_{c}^{\oplus f(c)}$ for all $f\in\operatorname{KRS}(C,e)$.
Remark 3.5.
For a tame partial KRS-monoid $X=(X,e,\oplus)$ we have $X\cong\operatorname{KRS}(X_{\mathrm{ind}},e^{\prime})$, where $e^{\prime}$ is the restriction of $e$ to $X_{\mathrm{ind}}\times X_{\mathrm{ind}}$. In this situation, any map $\mathbf{g}:C\rightarrow\mathbb{Z}^{n}$ can be extended to a framing $\mathbf{g}:\operatorname{KRS}(C,e)\rightarrow\mathbb{Z}^{n}$ by simply setting $\mathbf{g}(f):=\sum_{c\in C}f(c)\mathbf{g}(c)$.
3.2. KRS-monoids of $\tau^{-}$-reduced components
Let $A$ be a basic finite dimensional algebra over an algebraically closed field $𝕜\Bbbk$. We may identify the Grothendieck group $K_{0}(A)$ of $A$ with $\mathbb{Z}^{n}$ for some non-negative integer $n$. Consider the set $\operatorname{DecIrr}^{\tau^{-}}(A)$ of decorated, generically $\tau^{-}$-reduced, irreducible components of the representation varieties of $A$. See Section 1.3 for the terminology, and [6, Sec. 3 and 5] for more details. The set $\operatorname{DecIrr}^{\tau^{-}}(A)$ becomes a partial KRS-monoid under the function
$e_{A}:\operatorname{DecIrr}^{\tau^{-}}(A)\times\operatorname{DecIrr}^{\tau^{-}}(A)\rightarrow\mathbb{N}$
that to each pair $(X,Y)\in\operatorname{DecIrr}^{\tau}(A)\times\operatorname{DecIrr}^{\tau}(A)$ associates the generic value $e_{A}(X,Y)$ that on the Zariski product $X\times Y$ takes the symmetrized (injective) $E$-invariant
(3.1) $E_{A}((M,\mathbf{v}),(N,\mathbf{w})):=\\ \dim\operatorname{Hom}_{A}(\tau^{-1}M,N)+\dim\operatorname{Hom}_{A}(\tau^{-1}N,M)+\underline{\dim}(M)\cdot\mathbf{w}+\mathbf{v}\cdot\underline{\dim}(N);$
and the partially-defined sum
:$\oplus:\{(X,Y)\in\operatorname{DecIrr}^{\tau^{-}}(A)\times\operatorname{DecIrr}^{\tau^{-}}(A)\mid e(x,y)=0\}\rightarrow\operatorname{DecIrr}^{\tau^{-}}(A)$
given by the direct sum of $e_{A}$-orthogonal irreducible components, $X\oplus Y:=\overline{X\oplus Y}$.
The function $\mathbf{g}_{A}:\operatorname{DecIrr}^{\tau^{-}}(A)\rightarrow\mathbb{Z}^{n}$ that to each $X\in\operatorname{DecIrr}^{\tau}(A)$ associates the generic value taken on $X$ by the (injective) $g$-vector
$\mathbf{g}_{A}(M,\mathbf{v}):=(\dim\operatorname{Hom}(\tau^{-1}M,S_{i})-\dim(S_{i},M)+v_{i})_{i\in Q_{0}}$
is a framing for the partial KRS-monoid $(\operatorname{DecIrr}^{\tau^{-}}(A),e_{A},\oplus)$. More precisely:
Theorem 3.6.
Let $A$ a finite-dimensional basic algebra over an algebraically closed field. Let $n$ be the rank of the Grothendieck group of $A$.
  • (a)
    $\operatorname{DecIrr}^{\tau^{-}}(A)=(\operatorname{DecIrr}^{\tau^{-}}(A),e_{A},\oplus,\mathbf{g}_{A})$ is a framed, free KRS-monoid of rank $n$. The subset $\operatorname{DecIrr}^{\tau}_{\mathrm{ind}}(A)$ of components, which contain a dense set of indecomposable representations is precisely the set of indecomposable elements in the sense of Definition 3.3.
  • (b)
    With the framing from the generic (injective) $g$-vector, $\operatorname{DecIrr}^{\tau^{-}}(A)$ is a free partial KRS-monoid of rank $n$.
  • (c)
    If $A$ is of tame representation type, then $\operatorname{DecIrr}^{\tau^{-}}(A)$ is tame in the sense of Definition 3.3. Consequently, there is an isomorphism of partial KRS-monoids
    $\operatorname{DecIrr}^{\tau^{-}}(A)\cong\operatorname{KRS}(\operatorname{DecIrr}^{\tau^{-}}_{\mathrm{ind}}(A),e_{A}).$
    Moreover in this case each $Z\in\operatorname{DecIrr}^{\tau}_{\mathrm{ind}}(A)$ contains either a dense orbit, or a one-parameter family of bricks.
Remark 3.7.
  1. (1)
    In the form it is stated, Theorem 3.6 appeared (up to duality) in [24]. Its assertions are restatements of previous results by various authors: Part (a) is a well-known combination of [7, Theorem 1.2] and [6, Theorems 1.3 and 1.5]. Part (b) is a theorem of Plamondon [47]. Part (c) is [26, Corollary 1.7] and [23, Theorem 3.2].
  2. (2)
    More precisely, Theorem 3.6 is just the dual version of [24, Thm. 2.5]. There, the result was stated in terms of generically $\tau$-reduced components with the function $e_{A}^{\operatorname{proj}}$ which is defined in terms of the projective symmetrized E-invariant
    $\qquad\quad E_{A}^{\operatorname{proj}}((M,\mathbf{v}),(N,\mathbf{w})):=\\ \dim\operatorname{Hom}_{A}(M,\tau N)+\dim\operatorname{Hom}_{A}(N,\tau M)+\mathbf{v}\cdot\underline{\dim}(N)+\underline{\dim}(M)\cdot\mathbf{w},$
    with framing defined by the generic values of the projective $g$-vector
    $\mathbf{g}_{A}^{\operatorname{proj}}(M,\mathbf{v}):=(\dim\operatorname{Hom}_{A}(S_{i},\tau M,S_{i})-\dim\operatorname{Hom}_{A}(M,S_{i})+v_{i})_{i\in Q_{0}}$
    Duality over the ground field $𝕜\Bbbk$ induces a natural isomorphism of framed partial KRS-monoids
    $(\operatorname{DecIrr}^{\tau^{-}}(A),e_{A},\oplus,\mathbf{g}_{A})\cong(\operatorname{DecIrr}^{\tau}(A^{\operatorname{op}}),e^{\operatorname{proj}}_{A^{\operatorname{op}}},\oplus,\mathbf{g}^{\operatorname{proj}}).$
[…]
In view of [10, Proposition 10.4 and Remark 10.8], we have
(3.2) $\mathbf{g}_{A}(M)=(\dim\operatorname{ker}(M(\gamma_{k}))-\dim M(k))_{k\in Q_{0}}=:\mathbf{g}_{A}^{\mathrm{DWZ}}(M,\mathbf{0}).$
Here $\mathbf{g}_{A}^{\mathrm{DWZ}}(M)$ denotes the $g$-vector of a QP-representation in the sense of [10, (1.13)]. This can be easily extended to decorated QP-representations.
[…]
(3.3) $\gamma_{k}^{B}:\mathbb{Z}^{Q_{0}}\rightarrow\mathbb{Z}^{Q_{0}}\quad\text{with}\quad\gamma_{k}^{B}(\mathbf{g})_{i}=\begin{cases}-g_{i}&\text{if }i=k,\\ g_{i}+\text{sgn}(g_{k})[b_{ik}\cdot g_{k}]_{+}&\text{else},\end{cases}$
where $B(Q)=(b_{ij})$ is the matrix defined in (1.2). Note that, this is just another way of writing the transformation rule for g-vectors from [19, Conj. 7.12], which was proved in [10, Thm. 1.7]. Compare also with Equation (5.6).
Proposition 3.8.
[24, Proposition 2.8] For each $Z\in\operatorname{DecIrr}^{\tau^{-}}(A)$ there exists a dense open subset $U_{Z}\subset Z$, a unique irreducible component $\tilde{\mu}_{k}(Z)\in\operatorname{DecIrr}^{\tau^{-}}(A^{\prime})$ and a regular map $\nu_{Z}:U_{Z}\rightarrow\tilde{\mu}_{k}(Z)$ with the following properties:
  • (a)
    For each $X\in U_{Z}$ we have $\nu_{Z}(X)\cong\mu_{k}(X)$, where $\mu_{k}$ denotes the mutation of the decorated QP-representation $X$ in direction $k$, as defined in [9].
  • (b)
    The morphism of affine varieties $G_{\mathbf{d}(\tilde{\mu}_{k}(Z))}\times U_{Z}\rightarrow\tilde{\mu}_{k}(Z),(g,X)\mapsto g.\nu_{Z}(X)$ is dominant.
  • (c)
    For each $Z\in\operatorname{DecIrr}^{\tau}(A)$ we have
    $\mathbf{g}_{A^{\prime}}(\tilde{\mu}_{k}(Z))=\gamma_{k}^{B}(\mathbf{g}_{A}(Z)).$
  • (d)
    $\tilde{\mu}_{k}(\tilde{\mu_{k}}(Z))=Z$ for all $Z\in\operatorname{DecIrr}^{\tau^{-}}(A)$.
  • (e)
    The map
    $\tilde{\mu}_{k}:\operatorname{DecIrr}^{\tau^{-}}(A)\rightarrow\operatorname{DecIrr}^{\tau^{-}}(A^{\prime})$
    is an isomorphism of partial KRS-monoids. In particular, $Z\in\operatorname{DecIrr}^{\tau^{-}}(A)$ is indecomposable if and only if $\tilde{\mu}_{k}(Z)\in\operatorname{DecIrr}^{\tau^{-}}(A^{\prime})$ is indecomposable.
Remark 3.9.
(1) Similar to Theorem 3.6, the above Proposition 3.8 is the dual version of the original one in [24], see also the discussion in Remark 3.7.
(2) In view of the definition of the piecewise linear map $\gamma_{k}^{B}$ in Equation (3.3) and part (c) of Proposition 3.8 above we have the following interesting equality
$\left[\begin{array}[]{c}-\mu_{k}(B(Q))\\ \phantom{-}\mathbf{g}_{A}^{\prime}(\tilde{\mu}_{k}(Z))\end{array}\right]=\mu_{k}\left(\left[\begin{array}[]{c}-B(Q)\\ \phantom{-}\mathbf{g}_{A}(Z)\end{array}\right]\right)$
of $(n+1)\times n$ matrices, where $\mu_{k}$ is Fomin-Zelevinsky’s matrix mutation. Thus, Proposition 3.8 is a generalization of Nakanishi-Zelevinsky’s result [45, Equation (4.1)].
3.3. KRS-monoids of laminations
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a possibly-punctured surface with marked points. In [24, Section 4], we considered the set $\mathcal{LC}^{*}(\mathbf{\Sigma})$ of homotopy classes of marked curves and loops on $\mathbf{\Sigma}$ that have no kinks. Inspired by [50], we introduce a symmetric marked intersection number
$\operatorname{Int}^{*}_{\mathbf{\Sigma}}:\mathcal{LC}^{*}(\mathbf{\Sigma})\times\mathcal{LC}^{*}(\mathbf{\Sigma})\rightarrow\mathbb{Z}_{\geq 0}.$
Given a marked curve (or loop) $(\gamma,c)\in\mathcal{LC}^{*}(\mathbf{\Sigma})$, let $(\gamma,c)^{-1}$ denote the inversely oriented curve $\gamma$, with accordingly swapped decoration $c$. The involution $(\gamma,c)\mapsto(\gamma,c)^{-1}$ induces an equivalence relation $\simeq$ on $\mathcal{LC}^{*}(\mathbf{\Sigma})$. For us, as in [50], a simple marked curve is a marked curve $(\gamma,c)\in\mathcal{LC}^{*}(\mathbf{\Sigma})$ which has self-intersection number 00. We let $\mathcal{LC}^{*}_{\tau}(\mathbf{\Sigma})$ denote the set of all simple marked curves, and define the tame partial KRS-monoid of laminations
$\operatorname{Lam}(\mathbf{\Sigma}):=\operatorname{KRS}(\mathcal{LC}^{*}_{\tau}(\mathbf{\Sigma})/_{\simeq},\operatorname{Int}^{*}_{\mathbf{\Sigma}}).$
See Example 3.4 and [24, Section 4.3] for more details. The set $\mathcal{LC}^{*}_{\tau}(\mathbf{\Sigma})/_{\simeq}$ can be identified with the set of laminations $\mathcal{C}^{\circ}(\mathbf{\Sigma})$ considered in [44] (two marked curves having intersection number equal to 00 corresponds to them being $\mathcal{C}^{\circ}$-compatible in Musiker-Schiffler-Williams’ nomenclature). This is also compatible with the treatment of laminations in [16].
The tame partial KRS-monoid $\operatorname{Lam}(\mathbf{\Sigma})$ can be equipped with a framing coming from (dual) shear coordinates. We shall describe this framing only for the indecomposable elements of $\operatorname{Lam}(\mathbf{\Sigma})$, which we call laminates. For the reader’s convenience, we fully recall from [24] the definition of the vector of dual shear coordinates of a laminate with respect to an arbitrary tagged triangulation. Our treatment essentially follows [16], but with slight changes in conventions –hence the adjective dual in the term dual shear coordinates. Let $T$ be a tagged triangulation of $(\Sigma,\mathbb{M},\mathbb{P})$, and let $\lambda$ be either a tagged arc or a simple closed curve on $(\Sigma,\mathbb{M},\mathbb{P})$. We define the vector $\operatorname{Sh}_{T}(\lambda)=(\operatorname{Sh}_{T}(\lambda)_{\gamma})_{\gamma\in T}$ of dual shear coordinates of $\lambda$ with respect to $T$ in steps as follows.
Case 1.
Assume that $T$ has non-negative signature $\delta_{T}:\mathbb{P}\rightarrow\{1,0,-1\}$. Then, at any given puncture $p$, either all tagged arcs in $T$ incident to $p$ are tagged plain, or exactly two tagged arcs in $T$ are incident to $p$, their underlying ordinary arcs being isotopic to each other, their tags at $p$ differing from one another, and the tags at their other endpoint being plain.
Following [15], represent $T$ through an ideal triangulation $T^{\circ}$ defined as follows. For each $p\in\mathbb{P}$ such that $\delta_{T}(p)=0$, let $i_{p},j_{p},$ be the two tagged arcs in $T$ that are incident to $p$, with $i_{p}$ tagged plain at $p$, and $j_{p}$ tagged notched at $p$. Then $T^{\circ}:=\{\gamma^{\circ}\ |\ \gamma\in T\}$ is obtained from $T$ by setting $\gamma^{\circ}:=\gamma$ for every $\gamma\in T$ both of whose ends are tagged plain, and by replacing each $j_{p}$ with a loop $j_{p}^{\circ}$ closely surrounding $i_{p}^{\circ}$ for $i$ and $j$ as above. Thus, the corresponding ordinary arcs $i_{p}^{\circ},j_{p}^{\circ}\in T^{\circ}$ form a self folded triangle that has $i_{p}^{\circ}$ as the folded side, and $j_{p}^{\circ}$ as the loop closely surrounding $i_{p}^{\circ}$. See [15, Sections 9.1 and 9.2].
If $\lambda$ is a simple closed curve, set $L:=\lambda$. Otherwise, let $L$ be the curve obtained from $\lambda$ by modifying its two ending segments according to the following rules:
  • any endpoint incident to a marked point in the boundary is slightly slided along the boundary segment lying immediately to its right as in Figure 1 (here, we stand upon the surface using its orientation, and look from the marked point towards the interior of surface, note that we use the orientation of $\Sigma$ to determine what is right and what is left);
    Figure 1. Slightly sliding endpoints lying on the boundary
  • any ending segment of $\lambda$ tagged plain at a puncture $q$ is replaced with a non-compact curve segment spiralling towards $q$ in the clockwise sense;
  • any ending segment of $\lambda$ tagged plain at a puncture $q$ is replaced with a non-compact curve segment spiralling towards $q$ in the counter-clockwise sense.
Take an arc $\gamma\in T$. In order to define the shear coordinate $\operatorname{Sh}_{T}(\lambda)_{\gamma}$ we need to consider two subcases.
Subcase 1.1.
Suppose that the ordinary arc $\gamma^{\circ}\in T^{\circ}$ is not the folded side of a self-folded triangle of $T^{\circ}$. Then $\gamma^{\circ}$ is contained in exactly two ideal triangles of $T^{\circ}$, and the union ¯$\overline{\lozenge}_{\gamma^{\circ}}$ of these two triangles is either a quadrilateral (if $\gamma^{\circ}$ does not enclose a self-folded triangle) or a digon (if $\gamma^{\circ}$ encloses a self-folded triangle). In any of these two situations, the complement in ¯$\overline{\lozenge}_{\gamma^{\circ}}$ of the union of the arcs belonging to $T^{\circ}\setminus\{\gamma^{\circ}\}$ can be thought to be an open quadrilateral $\lozenge_{\gamma^{\circ}}$ in which $\gamma^{\circ}$ sits as a diagonal. The shear coordinate $\operatorname{Sh}_{T}(\lambda)_{\gamma}$ is defined to be the number of segments of $\lozenge_{\gamma^{\circ}}\cap L$ that form the shape of a letter $Z$ when crossing $\gamma^{\circ}$ minus the number of segments of $\lozenge_{\gamma^{\circ}}\cap L$ that form the shape of a letter $S$ when crossing $\gamma^{\circ}$.
Subcase 1.2.
Suppose that the ordinary arc $\gamma^{\circ}\in T^{\circ}$ is the folded side of a self-folded triangle of $T^{\circ}$. Then there is a puncture $p\in\mathbb{P}$ such that $\gamma=i_{p}$ and $\gamma^{\circ}=i_{p}^{\circ}$. Define
$\operatorname{Sh}_{T}(\lambda)_{i_{p}}:=\operatorname{Sh}_{T}(\lambda^{\prime})_{j_{p}}.$
where $\lambda^{\prime}$ is obtained from $\lambda$ by switching the tags of $\lambda$ at the puncture $p$.
Case 2.
Assume now that $T$ is an arbitrary tagged triangulation of $(\Sigma,\mathbb{M},\mathbb{P})$. The set of punctures at which $T$ has negative signature is the inverse image $\delta_{T}^{-1}(-1)$. Set $T^{\prime}$ to be the tagged triangulation obtained from $T$ by changing from notched to plain all the tags incident to punctures in $\delta_{T}^{-1}(-1)$. Thus, $T^{\prime}$ is a tagged triangulation of signature zero, so dual shear coordinates with respect to $T^{\prime}$ have already been defined. Set
$\operatorname{Sh}_{T}(\lambda):=\operatorname{Sh}_{T^{\prime}}(\lambda^{\prime}),$
where $\lambda^{\prime}$ is obtained from $\lambda$ by switching all the tags of $\lambda$ at the punctures belonging to the set $\delta_{T}^{-1}(-1)$.
Example 3.10.
In Figure 2,
Figure 2. Left: computation of the vector of dual shear coordinates $\operatorname{Sh}_{T}(\lambda)$. Right: computation of the vector of Fomin–Thurston’s shear coordinates $\operatorname{Sh}_{\overline{T}}^{\operatorname{FST}}(\overline{\lambda})$.
the reader can glimpse the relation between the dual shear coordinates we have defined above and the shear coordinates used by Fomin–Thurston, namely,
(3.4) $\operatorname{Sh}_{T}(\lambda)=\operatorname{Sh}_{\overline{T}}^{\operatorname{FST}}(\overline{\lambda}),$
where $\operatorname{Sh}_{T}(\lambda)$ is the vector of dual shear coordinates we have defined above, $\overline{T}$ and $\overline{\lambda}$ are the images of $T$ and $\lambda$ in the surface obtained as the mirror image of $(\Sigma,\mathbb{M},\mathbb{P})$, and $\operatorname{Sh}_{\overline{T}}^{\operatorname{FST}}(\overline{\lambda})$ is the vector of shear coordinates of Fomin–Thurston, [16, Definition 13.1].
With these definitions in place we can restate [16, Theorem 13.5] for dual shear coordinates as follows.
Theorem 3.11.
Let $T$ and $T^{\prime}$ be tagged triangulations of $(\Sigma,\mathbb{M},\mathbb{P})$ related by the flip of a tagged arc $k\in T$, and let $\lambda$ be either a tagged arc or a simple closed curve on $(\Sigma,\mathbb{M},\mathbb{P})$. Then
$\left[\begin{array}[]{c}-B(T^{\prime})\\ \operatorname{Sh}_{T^{\prime}}(\lambda)\end{array}\right]=\mu_{k}\left(\left[\begin{array}[]{c}-B(T)\\ \operatorname{Sh}_{T}(\lambda)\end{array}\right]\right).$
In view of the definition of the piecewise linear map $\gamma_{k}^{B}$ in Equation (3.3) and the definition of the Fomin-Zelevinsky matrix mutation rule, Theorem 3.11 implies in particular
(3.5) $\operatorname{Sh}_{T^{\prime}}(\lambda)=\gamma^{B(T)}_{k}(\operatorname{Sh}_{T}(\lambda)).$
See also Remark 3.9 (2).
3.4. The isomorphism of framed KRS-monoids $\operatorname{Lam}(\mathbf{\Sigma})\cong\operatorname{DecIrr}^{\tau}(A)$
Generalizing [23, Theorem 10.13 and Proposition 10.14] from the unpunctured to the punctured case, the main result of [24] states with our conventions
Theorem 3.12.
[24, Theorem 1.1.] Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with non-empty boundary. For each tagged triangulation $T$ of $\mathbf{\Sigma}$ there is a unique isomorphism of partial KRS-monoids
$\pi_{T}:(\operatorname{Lam}(\mathbf{\Sigma}),\operatorname{Int}^{*},+)\rightarrow(\operatorname{DecIrr}^{\tau-}(A(T)),e_{A(T)},\oplus),$
such that the diagram of functions and sets
[…]
where $A(T)$ is the dual Jacobian algebra of the quiver with potential associated to $T$ in [5, Definition 4.1], [36, Definitions 3.1 and 3.2], and $\mathbf{g}_{A(T)}$ is the generic injective g-vector.
Roughly speaking, the proof strategy followed in [24] consists of two steps:
  1. (1)
    the statement of the theorem is shown to hold for tagged triangulations of signature zero, exploiting heavily the fact that for those triangulations the Jacobian algebra is skewed-gentle, and that in this situation a combinatorial description of the generically $\tau^{-}$-reduced components is now available [20];
  2. (2)
    given a tagged triangulation $T$ for which the theorem holds, and another tagged triangulation $T^{\prime}$ obtained by flipping an arc $k\in T$, Proposition 3.8 and Theorem 3.11 are applied to produce a commutative diagram
[…]
    and this allows to deduce that the theorem holds then for $T^{\prime}$ as well.
Since any two tagged triangulations of a surface with non-empty boundary are related by a finite sequence of flips, and since any such surface certainly admits at least one triangulation of signature zero, these two steps yield a proof of Theorem 3.12.
4. Triangulations adapted to closed curves
Let $\alpha$ be an essential loop on $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$, i.e., a non-contractible simple closed curve that is furthermore not contractible to a puncture, and let $T=\{\tau_{1},\ldots,\tau_{n}\}$ be a tagged triangulation of $\mathbf{\Sigma}$. The band graph associated to $\alpha$ with respect to $T$, cf. [44, Definition 3.4 and §8.3], will be denoted $G(T,\alpha)$. The polynomial $F_{\zeta}^{T}$ defined in [44, Definition 3.14] (see also [44, §8.3]), with $\zeta:=\alpha$, will be called snake $F$-polynomial and denoted $F_{G(T,\alpha)}$. Similarly, the integer vector from [44, Definition 6.1] will be called the snake $g$-vector and denoted $\mathbf{g}_{G(T,\alpha)}$. Notice that
$F_{G(T,\alpha)}=\displaystyle\sum_{P\in\mathcal{P}(G(T,\alpha))}y(P)$
where $\mathcal{P}(G(T,\alpha))$ is the collection of all good matchings of the band graph $G(T,\alpha)$. Furthermore, by [44, Proposition 6.2], we have
$\mathbf{g}_{G(T,\alpha)}=\deg\left(\frac{x(P_{-})}{\operatorname{cross}(T,\alpha)}\right),$
where $P_{-}$ is the minimal matching of the band graph $G(T,\alpha)$, cf. [44, Definition 3.7], and $\deg:\mathbb{Z}[x_{1}^{\pm 1},\ldots,x_{n}^{\pm 1},y_{1},\ldots,y_{n}]\rightarrow\mathbb{Z}^{n}$ is the grading defined by
$\deg(x_{i})=\mathbf{e}_{i}\qquad\text{and}\qquad\deg(y_{j}):=-\mathbf{b}(T)_{j}$
as in [19, Proposition 6.1], $\mathbf{b}(T)_{j}$ being the $j^{\operatorname{th}}$ column of $B(T)$, see Definition 2.9.
Proposition 4.1.
For each non-contractible simple closed curve $\alpha$ on $\mathbf{\Sigma}$ there exists a triangulation $T_{\alpha}$ without self-folded triangles such that:
  • the Jacobian algebra of the restriction $(Q^{\prime}(T_{\alpha})|_{I},{S^{\prime}}(T_{\alpha})|_{I})$ is gentle, where $I$ is the set of arcs crossed by $\alpha$;
  • the band module $M(T_{\alpha},\alpha,\lambda,1)$ over $\Lambda_{I}:=\mathcal{P}(Q(T_{\alpha})|_{I},S(T_{\alpha})|_{I})$ of quasi-length $1$ associated to the parameter $\lambda\in\mathbb{C}^{*}$ is a module over the Jacobian algebra $A(T_{\alpha}):=\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T_{\alpha}),S^{\prime}(T_{\alpha}))$ as well and satisfies:
    $F_{M(T_{\alpha},\alpha,\lambda,1)}$ $=F_{G(T,\alpha)}$ $E_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))$ $=1$
    $\mathbf{g}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))$ $=\operatorname{Sh}_{T_{\alpha}}(\alpha)=\mathbf{g}_{T_{\alpha}}(G(T_{\alpha},\alpha)).$
Proof.
Cutting $\mathbf{\Sigma}$ along $\alpha$ we are left with at most two marked surfaces $(S_{1},M_{1})$ and $(S_{2},M_{2})$ – note that if $\alpha$ is non-separating then one $S_{i}$ will be empty. To construct $T_{\alpha}$ we split the task into two cases.
Case 1.
If $M_{i}\neq\emptyset$ for any $i\in\{1,2\}$ then there exist compatible arcs $\gamma_{1}$ and $\gamma_{2}$ in $\mathbf{\Sigma}$ which bound $\alpha$ in a cylinder $C_{1,1}$. Define $T_{\alpha}$ to be any triangulation containing both $\gamma_{1}$ and $\gamma_{2}$, with the extra property of not having self-folded triangles. See Figure 3.
Figure 3. The two types of $\alpha$ occurring in Case 1, together with the arcs $\gamma_{1}$ and $\gamma_{2}$ bounding it in a cylinder.
Case 2.
Otherwise, $\alpha$ is a separating curve where $S_{i}\neq\emptyset$ and $M_{i}=\emptyset$ for some $i\in\{1,2\}$. In this case there exists an arc $\gamma\in\mathbf{\Sigma}$ bounding $\alpha$ in a bordered surface $\mathbf{\Sigma}_{g,1}$ of genus $g$ with one boundary component and one marked point. Define $T_{\alpha}$ to be any triangulation containing $\gamma$, with the extra property of not having self-folded triangles. See Figures 4 and 6.
Figure 4. The type of $\alpha$ occurring in Case 2, together with the arc $\gamma$ bounding it in a genus $g$ surface with one boundary component and one marked point.
In both cases, it is clear that $\Lambda_{I}:=\mathcal{P}(Q(T_{\alpha})|_{I},S(T_{\alpha})|_{I})$ is a gentle algebra. That $M(T_{\alpha},\alpha,\lambda,1)$ is a module also over $A(T_{\alpha})$ follows immediately from [9, Proposition 8.9] or by direct inspection.
It is easy to check that $A(\overline{T_{\alpha}})=A(T_{\alpha})^{\operatorname{op}}$ and $M(\overline{T_{\alpha}},\overline{\alpha},\lambda,1)\cong D(M(T_{\alpha},\alpha,\lambda,1))$ as representations of $A(\overline{T_{\alpha}})$. Therefore, by (3.2), [34, proof of Theorem 10.0.5] and (3.4), we have
$\mathbf{g}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))=\mathbf{g}_{A(\overline{T_{\alpha}})}^{{\operatorname{proj}}}(M(\overline{T_{\alpha}},\overline{\alpha},\lambda,1))=\operatorname{Sh}_{\overline{T_{\alpha}}}^{\operatorname{FST}}(\overline{\alpha})=\operatorname{Sh}_{T_{\alpha}}(\alpha).$
The equality $\operatorname{Sh}_{T_{\alpha}}(\alpha)=\mathbf{g}_{G(T_{\alpha},\alpha)}$ follows from [23, Proposition 10.14].
Recall, that our quiver $Q^{\prime}(T_{\alpha})$ is opposite to the quiver $Q(T)$ which was used in [34] and in [23].
Lemma 4.2.
The band graph $G(T_{\alpha},\alpha)$ is a ‘zig-zag’ band graph whose corresponding band is given in Figure 5.
Figure 5. Here we depict the possible shapes of what we call ‘zig-zag’ band graphs. The only restriction is that $n\geq 2$.
Proof.
When $T_{\alpha}$ is defined via case $1$, then a direct computation obtains the $n=2$ band graph in Figure 5.
So, suppose $T_{\alpha}$ is defined via case $2$. Then there exists a unique triangle $\Delta$ of $T_{\alpha}$, lying inside $\mathbf{\Sigma}_{g,1}$, such that $\gamma$ is an edge of $\Delta$. Let us denote the remaining two edges of $\Delta$ by $\gamma_{1}$ and $\gamma_{2}$. For some $\beta$ and $\delta$ in $T$ we see that $\alpha$ has the following sequences of ‘zig-zag’ intersections: $\gamma_{1},\gamma_{2},\beta$ and $\delta,\gamma_{1},\gamma_{2}$. All other sequences of intersections (of $\alpha$ with $T$) of length $3$ are ‘fan’ intersections. This completes the proof. ∎
Figure 6. The green shaded area denotes the triangle $\Delta$ used in the proof of Lemma 4.2.
Lemma 4.3.
For $\lambda\in\mathbb{C}^{*}$, the endomorphism ring $\operatorname{End}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))$ of the regular quasi-simple band module $M(T_{\alpha},\alpha,\lambda,1)$ is isomorphic to $\mathbb{C}$.
Proof.
The band graph of $\alpha$ with respect to $T_{\alpha}$ is of ‘zig-zag’ type by construction. As shown in Lemma 4.2, each ‘zig-zag’ band graph gives rise to a band $b$ of the form shown in (4.1). Moreover, the ”support” $\Lambda_{I}$ of $b$ is a special biserial (in fact hereditary) algebra. Thus we have trivially $\operatorname{End}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))=\operatorname{End}_{\Lambda_{I}}(M(T_{\alpha},\alpha,\lambda,1))$ We shall follow the setup and terminology used by Butler and Ringel in [1] for representations of band modules.
Let $M=M(b,\lambda,1)$ be a regular simple band module corresponding to the following band $b$:
[…]
Here, we may think of vertices $t_{i_{j}}\in Q_{0}(T_{\alpha})$ as the arcs of the triangulation $T_{\alpha}$ which intersect consecutively with the loop $\alpha$. Note that possibly $t_{i_{j}}=t_{i_{k}}$ for certain $i\neq k$. However, $t_{i_{1}}\neq t_{i_{n}}$ since the arrow $\alpha_{n}$ is not a loop. Specifically, $M$ is defined by the following collection of vector spaces and maps:
${\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$${\ldots}{\mathbb{C}}$${\mathbb{C}}$${\mathbb{C}}$$I$$I$$I$$I$$\lambda^{-1}$
Now, we can use Krause’s formalism from [33]. With the notation from loc. cit. p. 191 we denote by $\mathcal{A}(b,b)$ the set of isomorphism classes of admissable (sic) triples for the pair of bands $(b,b)$. Since the band $b$ has a unique source, which is mapped to $t_{i_{1}}$ and a unique sink, which is mapped to $t_{i_{n}}\neq t_{i_{1}}$ we have $\mathcal{A}(b,b)=\{[(b,\operatorname{id}_{b},\operatorname{id}_{b})]\}$. Thus, $\operatorname{End}_{\Lambda_{I}}(M)\cong\mathbb{C}$ follows directly from the Theorem on [33, p. 191]. ∎
By Lemma 4.3 we have $\operatorname{End}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))=\mathbb{C}$. On the other hand, one can directly check that $\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))\cdot\mathbf{g}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))=0$. Hence
$E_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))$ $=\dim(\operatorname{End}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1)))+$
$\quad+\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))\cdot\mathbf{g}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))=1.$
The equality $F_{M(T_{\alpha},\alpha,\lambda,1)}=F_{G(T_{\alpha},\alpha)}$ follows from [23, Lemma 11.4 and Remark 11.7] and [29, Theorem 1.2] as in the proof of [23, Proposition 11.8]. Note however, that due to our convention for $A(T_{\alpha})$ resp. for $\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T_{\alpha})\mid_{I},S^{\prime}(T_{\alpha})_{I})$, we can avoid the use of dual CC-functions. This finishes the proof of Proposition 4.1. ∎
Remark 4.4.
In the proof of Proposition 4.1 we have partially used the explicit definition of the potential $S(T)$: we know that the restriction of $S(T)$ to the set $I$ is a gentle algebra.
Corollary 4.5.
In the situation of Proposition 4.1,
  • the Zariski closure
    $Z_{T_{\alpha},\alpha}:=\overline{\bigcup_{\lambda\in\mathbb{C}^{*}}\operatorname{GL}_{\mathbf{d}}(\mathbb{C})\cdot M(T_{\alpha},\alpha,\lambda,1)}$
    is a generically $\tau^{-}$-reduced indecomposable irreducible component of the representation space $\operatorname{rep}(A(T_{\alpha}),\mathbf{d})$, where $\mathbf{d}:=\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))$;
  • the generic (injective) $g$-vector $\mathbf{g}_{A(T_{\alpha})}(Z_{T_{\alpha},\alpha})$ is equal to the snake $g$-vector $\mathbf{g}_{G(T_{\alpha},\alpha)}$;
  • the generic value $CC_{A(T_{\alpha})}(Z_{T_{\alpha},\alpha})$ is equal to Musiker-Schiffler-Williams’ expansion $\operatorname{MSW}(G(T_{\alpha},\alpha))$.
Proof.
By [21, Theorem 7.1] (see also [35, Theorem 3.5]), the algebra $A(T_{\alpha})$ is tame. Furthermore, for $\lambda_{1}\neq\lambda_{2}$, the band modules $M(T_{\alpha},\alpha,\lambda_{1},1)$ and $M(T_{\alpha},\alpha,\lambda_{2},1)$ are non-isomorphic, and by Proposition 4.1, for every point $M$ in the dense open subset $\bigcup_{\lambda\in\mathbb{C}\setminus\{0\}}\operatorname{GL}_{\mathbf{d}}(\mathbb{C})\cdot M(T_{\alpha},\alpha,\lambda,1)$ we have $E_{A(T_{\alpha})}(M)=1$. Hence $Z_{T_{\alpha},\alpha}$ is a generically $\tau^{-}$-reduced indecomposable irreducible component by [23, Lemma 3.1 and Theorem 3.2] (see also [3, Section 2.2]).
By Proposition 4.1, for every point $M\in\bigcup_{\lambda\in\mathbb{C}\setminus\{0\}}\operatorname{GL}_{\mathbf{d}}(\mathbb{C})\cdot M(T_{\alpha},\alpha,\lambda,1)$ we have
$F_{M}=F_{G(T_{\alpha},\alpha)}\qquad\text{and}\qquad\mathbf{g}_{A(T_{\alpha})}(M)=\mathbf{g}_{G(T_{\alpha},\alpha)}.$
This implies the second and third assertions. ∎
5. The Combinatorial Key Lemma
5.1. Derksen–Weyman–Zelevinsky’s representation-theoretic Key Lemma
The following result is the Key Lemma of Derksen, Weyman and Zelevinsky.
Theorem 5.1 (Lemma 5.2, [10]).
Let $\mathcal{M}$ be a finite-dimensional decorated representation of a non-degenerate quiver with potential $(Q,S)$, and define $\overline{\mathcal{M}}$ to be the mutation of $\mathcal{M}$ in direction $k\in\{1,\ldots,n\}$. Then
(5.1) $(y_{k}+1)^{h_{k}}F_{\mathcal{M}}(y_{1},\ldots,y_{n})=(y_{k}^{\prime}+1)^{h_{k}^{\prime}}F_{\overline{\mathcal{M}}}(y_{1}^{\prime},\ldots,y_{n}^{\prime}),$
where
  • $(\mathbf{y}^{\prime},B^{\prime})\in\mathbb{Q}(y_{1},\ldots,y_{n})$ is obtained from $(\mathbf{y},B_{\operatorname{DWZ}}(Q))$ by $Y$-seed mutation at $k$,
  • $h_{k}$ and $h_{k}^{\prime}$ are the $k^{th}$ components of the $\mathbf{h}$-vectors $\mathbf{h}_{\mathcal{M}}$ and $\mathbf{h}_{\overline{\mathcal{M}}}$, respectively, defined by [10, Equations (1.8) and (3.2)].
Moreover, the $\mathbf{g}$-vector $\mathbf{g}_{\mathcal{P}(Q,S)}^{\operatorname{DWZ}}(\mathcal{M})=(g_{1},\ldots,g_{n})$ defined by [10, Equation (1.13)] satisfies
(5.2) $g_{k}=h_{k}-h_{k}^{\prime}$
and is related to the $\mathbf{g}$-vector $\mathbf{g}_{\mathcal{P}(\mu_{k}(Q,S))}^{\operatorname{DWZ}}(\overline{\mathcal{M}})=(g_{1}^{\prime},\ldots,g_{n}^{\prime})$ via
(5.3) $g_{j}^{\prime}:=\left\{\begin{array}[]{ll}-g_{k},&\text{if $j=k$};\\ \\ g_{j}+[b_{jk}]_{+}g_{k}-b_{jk}h_{k},&\text{if $j\neq k$}.\\ \end{array}\right.$
Remark 5.2.
By [10, Proposition 10.4 and Remark 10.8], $\mathbf{g}_{\mathcal{M}}^{\operatorname{DWZ}}=\mathbf{g}_{\mathcal{M}}^{\operatorname{inj}}$ (see also (3.2)).
5.2. Statement of the Combinatorial Key Lemma
Theorem 5.3 (Combinatorial Key Lemma).
Let $T=\{\tau_{1},\ldots,\tau_{n}\}$ be a tagged triangulation of $\mathbf{\Sigma}$, $\tau_{k}$ a tagged arc belonging to $T$, and $\alpha$ a simple closed curve. We set $T^{\prime}$ to be the tagged triangulation obtained from $T$ by flipping $\tau_{k}$. The following equality holds:
(5.4) $(y_{k}+1)^{h_{k}}F_{G(T,\alpha)}(y_{1},\ldots,y_{n})=(y_{k}^{\prime}+1)^{h_{k}^{\prime}}F_{G(T^{\prime},\alpha)}(y_{1}^{\prime},\ldots,y_{n}^{\prime})$
where
  • $(\mathbf{y}^{\prime},B^{\prime})\in\mathbb{Q}(y_{1},\ldots,y_{n})$ is obtained from $(\mathbf{y},B(T))$ by mutation at $k$,
  • the snake $\mathbf{h}$-vector $\mathbf{h}_{G(T,\alpha)}$ is defined by setting
    $u^{h_{i}}:={F_{G(T,\alpha)}}_{{|}_{\operatorname{Trop}(u)}}(u^{[-b_{i1}]_{+}},\ldots,u^{-1},\ldots,u^{[-b_{in}]_{+}}),$
    where $u^{-1}$ is in the $i^{th}$ component. The snake $\mathbf{h}$-vector $\mathbf{h}_{G(T^{\prime},\alpha)}$ is defined analogously with respect to $T^{\prime}$ and $\mu_{k}(B(T))=B(T^{\prime})$.
Moreover, the snake $\mathbf{g}$-vector $\mathbf{g}_{G(T,\alpha)}=(g_{1},\ldots,g_{n})$ satisfies
(5.5) $g_{k}=h_{k}-h_{k}^{\prime}$
and $\mathbf{g}_{G(T,\alpha)}=(g_{1},\ldots,g_{n})$ is related to $\mathbf{g}_{G(T^{\prime},\alpha)}=(g_{1}^{\prime},\ldots,g_{n}^{\prime})$ by the following rule:
(5.6) $g_{j}^{\prime}:=\left\{\begin{array}[]{ll}-g_{k},&\text{if $j=k$};\\ \\ g_{j}+[b_{jk}]_{+}g_{k}-b_{jk}h_{k},&\text{if $j\neq k$}.\\ \end{array}\right.$
Remark 5.4.
Since the band graph $G(T,\alpha)$ always admits a minimal matching, the snake $F$-polynomial $F_{G(T,\alpha)}$ has constant term $1$, hence every entry of the snake $h$-vector $\mathbf{h}_{G(T,\alpha)}$ is non-positive.
Example 5.5.
As shown in Figure 7, let $T$ and $T^{\prime}$ be triangulations of the cylinder $C_{1,1}$, where $T^{\prime}$ is obtained from $T$ by the flip of $\tau_{1}$. Furthermore, let $\alpha$ be the unique simple closed curve of $C_{1,1}$.
Directly from the definition, let us compute the snake $F$-polynomials $F_{G(T,\alpha)}$ and $F_{G(T^{\prime},\alpha)}$ of $\alpha$. Namely, considering the good matchings associated to the band graphs $G(T,\alpha)$ and $G(T^{\prime},\alpha)$ shown in Figure 7 we get:
$F_{G(T,\alpha)}(y_{1},y_{2})=1+y_{2}+y_{1}y_{2}\hskip 28.45274pt\text{and}\hskip 28.45274ptF_{G(T^{\prime},\alpha)}(y_{1},y_{2})=1+y_{1}+y_{1}y_{2}.$
Now we wish to calculate the components $h_{1}$ and $h_{1}^{\prime}$ corresponding to $\tau_{1}$ in the respective $\mathbf{h}$-vectors $\mathbf{h}_{\alpha,T}$ and $\mathbf{h}_{\alpha,T^{\prime}}$. Since $b_{12}=-2$ and $b_{12}^{\prime}=2$ we get:
${F_{G(T,\alpha)}}_{|_{Trop(u)}}(u^{-1},u^{2})=1\oplus u^{2}\oplus u=1$
and
${F_{G(T^{\prime},\alpha)}}_{|_{\operatorname{Trop}(u)}}(u^{-1},1)=1\oplus u^{-1}\oplus u^{-1}=u^{-1}.$
Consequently, we have $h_{1}=0$ and $h_{1}^{\prime}=-1$. As such, (5.4) now reduces to:
$F_{G(T,\alpha)}(y_{1},y_{2})=(y_{1}^{\prime}+1)^{-1}F_{G(T^{\prime},\alpha)}(y_{1}^{\prime},y_{2}^{\prime}),$
which follows directly from the definition of $Y$-seed mutation. Indeed, $y_{1}^{\prime}:=y_{1}^{-1}$ and $y_{2}^{\prime}:=y_{2}(1+y_{1})^{2}$.
Figure 7. Left: triangulations $T$ and $T^{\prime}$ of $C_{1,1}$ related by the flip of $\tau_{1}$. Right: the band graphs $G(T,\alpha)$ and $G(T^{\prime},\alpha)$.
The rest of Section is devoted to proving Theorem 5.3.
5.3. Local segments and their local flips
Note that, in general, the band graphs $G(T,\alpha)$ and $G(T^{\prime},\alpha)$ will have very different shapes, consequently, their collections of good matchings seem to differ considerably. This is due to the fact that $\alpha$ may intersect the flip region multiple times, and in many different combinatorial ways. Nevertheless, we shall show that the relationships between the snake $\mathbf{g}$-vectors, snake $\mathbf{h}$-vectors and snake $F$-polynomials of $G(T,\alpha)$ and $G(T^{\prime},\alpha)$ are governed by ‘local’ considerations.
Let us (cyclically) label the intersections points between $\alpha$ and $T^{\circ}$ by $p_{1},p_{2},\ldots,p_{d}$. By convention, indices of these intersection points will always be taken modulo $d$.
Definition 5.6.
Consider a subcurve $\alpha_{ij}$ of $\alpha$ with intersection points $p_{i},\ldots,p_{j}$. We call $\alpha_{ij}$ a local segment of $\alpha$ with respect to $\gamma\in T$ if:
  • $\alpha_{ij}$ intersects $\gamma$ or $\gamma^{\prime}$
  • neither $p_{i-1}$ or $p_{j+1}$ lie on $\gamma$ or $\gamma^{\prime}$
  • $\alpha_{ij}$ is minimal with the above two properties (with respect to subcurve inclusion).
We list all possible local segments of $\alpha$, with respect to a tagged arc $\gamma\in T$, in Figure 8.
Figure 8. The complete list of local segments, considered up to rotations and reflections.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Remark 5.7.
Note that in all of the configurations listed in Figure 8, we have
$a_{1},a_{2},b_{1},b_{2},c_{1},c_{2}\notin\{\gamma,\gamma^{\prime}\}.$
Remark 5.8.
We have not included the case that $\alpha$ is a closed curve enclosed by $T$ in a cylinder $C_{1,1}$, but this is an exceptional case and can easily be checked separately.
We now list the snake graphs corresponding to the flips illustrated in Figure 8.
Figure 9. The list of snake graphs corresponding to Figure 8.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
As mentioned above, the strategy in proving Theorem 5.3 is to reduce the problem to checking that the theorem holds on the flips listed in Figure 8. To do this, we shall need a better understanding of the associated $\mathbf{h}$-vectors.
5.4. Local $\mathbf{h}$-vectors
Definition 5.9.
Let $\alpha_{ij}$ be a subcurve of $\alpha$ present in Figure 8. One can define a snake graph corresponding to the configuration, and thus obtain a local snake $F$-polynomial $F_{G(T,\alpha_{ij})}$. We call the $\mathbf{h}$-vector $\mathbf{h}_{G(T,\alpha_{ij})}$ defined through $F_{G(T,\alpha_{ij})}$ the local snake $\mathbf{h}$-vector of $\alpha$ with respect to $\alpha_{ij}$ and $T$.
The idea is to show that local snake $\mathbf{h}$-vectors naturally determine the whole snake $\mathbf{h}$-vector.
Proposition 5.10.
Let $\mathcal{L}$ be the set of all local segments of $\alpha$ with respect to $\gamma\in T$. Then for each $k\in\{1,\ldots,n\}$ we have the following equality:
$h_{\alpha;k}=\displaystyle\sum_{\alpha_{ij}\in\mathcal{L}}h_{\alpha_{ij};k}.$
To prove Proposition 5.10, we first show:
Lemma 5.11.
Keeping the terminology used in Proposition 5.10, the following inequality holds.
$\sum_{\alpha_{ij}\in\mathcal{L}}h_{\alpha_{ij};k}\leq h_{\alpha;k}.$
Proof.
Let $\gamma$ be any arc in $T$. Note that if $d$ is a diagonal of $G(T,\alpha)$ such that $b_{d\gamma}\neq 0$ then that diagonal appears as the diagonal of some tile in $G(T,\alpha_{ij})$. The lemma then follows by Remarks 5.4 and 5.7.
By Lemma 5.11, to prove Proposition 5.10 we need to show that there exists a good matching $P$ of $G(T,\alpha)$ such that the restriction of $P$ to any $G(T,\alpha_{ij})$ of $G(T,\alpha)$ achieves the corresponding local $h_{k}$. Namely, we shall introduce the idea of gluing together perfect matchings. With respect to this notion of gluing, the basic idea is show that for any $G(T,\alpha_{ij})$ and $G(T,\alpha_{lm})$, there exists perfect matchings $P_{ij}$ and $P_{lm}$, respectively, which can be glued together and which achieve their corresponding local $h_{k}$’s.
5.5. Extendability of perfect matchings
Definition 5.12.
Let $\mathcal{H}=[H_{1},\ldots,H_{s}]$ and $\mathcal{K}=[K_{1},\ldots,K_{t}]$ be snake graphs contained in a band graph $\mathcal{G}$, such that $H_{i}\neq K_{j}$ for all $i,j$. Let $P_{\mathcal{H}}$ and $P_{\mathcal{K}}$ be perfect matchings of $\mathcal{H}$ and $\mathcal{K}$, respectively. We say that $P_{\mathcal{H}}$ is extendable to $P_{\mathcal{K}}$ in $\mathcal{G}$ if there exists a perfect matching $P$ of the union $[H_{1},H_{2},\ldots,K_{t-1},K_{t}]$ such that the orientation induced by $P$ on the diagonals of $\mathcal{H}$ and $\mathcal{K}$ agrees with the orientation induced by $P_{\mathcal{H}}$ and $P_{\mathcal{K}}$, respectively.
Figure 10. An example and non-example of extendability. With respect to Definition 5.12, the tiles of $\mathcal{H}$ and $\mathcal{K}$ are shaded green and blue, respectively.
Definition 5.13.
Let $P$ be a perfect (resp. good) matching of a snake (resp. band) graph $\mathcal{G}$, and let $T$ be a tile of $\mathcal{G}$. We say that the diagonal of $T$ is positive with respect to $P$ if one of the following holds:
  • $T$ is an odd tile and the diagonal is oriented downwards;
  • $T$ is an even tile and the diagonal is oriented upwards.
Otherwise we say that the diagonal of $T$ is negative.
Lemma 5.14.
If $P_{\mathcal{H}}$ and $P_{\mathcal{K}}$ induce negative diagonals on the tiles $H_{s}$ and $K_{1}$, respectively, then $P_{\mathcal{H}}$ is extendable to $P_{\mathcal{K}}$.
Proof.
This follows from the existence of a minimal matching on $[H_{s},\ldots,K_{1}]$. ∎
Proof of Proposition 5.10.
Lemma 5.14 enables us to prove Proposition 5.10 with the following strategy. We shall show that for each $G(T,\alpha_{i})$ appearing in Figure 9 there exists a perfect matching $P_{i}$ such that:
  • $P_{i}$ achieves the corresponding local $h_{k}$;
  • $P_{i}$ induces a negative diagonal on the first and last tile of $G(T,\alpha_{i})$.
Case 1a): Here $b_{\gamma a}=-1$ and $b_{\gamma b}=1$. If the diagonal $\gamma$ is positive then so is the diagonal $a$. Therefore $h_{k}=0$. An alternative way to achieve $h_{k}=0$ is taking $P_{i}$ to be the minimal matching.
Case 1b): There is no diagonal labeled by $\gamma^{\prime}$ so $h_{k}\geq 0$. Taking $P_{i}$ to be the minimal matching achieves $h_{k}=0$.
Case 2a): Here $b_{\gamma a}=b_{\gamma c}=-1$. If the diagonal $\gamma$ is positive then so are the diagonals $a$ and $c$, hence $h_{k}=0$. So take $P_{i}$ to be the minimal matching.
Case 2b): Here $b_{\gamma^{\prime}a}=b_{\gamma^{\prime}c}=1$. Hence $h_{k}=-1$ is achieved by taking $P_{i}$ such that the diagonal $\gamma^{\prime}$ is positive, and all diagonals are negative.
Case 3a): Here $b_{\gamma a}=-2$ and $b_{\gamma b}=1$. If the diagonal $\gamma$ is positive then so is $a$, hence $h_{k}\geq 0$. Taking $P_{i}$ to be the minimal matching achieves $h_{k}=0$.
Case 3b): Here $b_{\gamma^{\prime}a}=2$ and $b_{\gamma^{\prime}b}=-1$. If the diagonal $\gamma^{\prime}$ is positive then so is the subsequent diagonal $b$, hence $h_{k}\geq 0$. Taking $P_{i}$ to be the minimal matching achieves $h_{k}=0$.
Case 4a): Here $b_{\gamma a}=-2$ and $b_{\gamma b}=b_{\gamma c}=1$. If any diagonal $\gamma$ is positive then so is the adjacent diagonal $a$, hence $h_{k}\geq 0$. Taking $P_{i}$ to be the minimal matching achieves $h_{k}=0$.
Case 4b): Here $b_{\gamma^{\prime}a}=2$ and $b_{\gamma^{\prime}b}=b_{\gamma^{\prime}c}=-1$. Taking $P_{i}$ such that diagonals $b$ and $c$ are negative, and all other diagonals are positive shows $h_{k}\leq 1-k$. Moreover, since there are only $k-1$ diagonals labelled by $\gamma^{\prime}$ then $h_{k}=1-k$.
Case 5a): There is no diagonal labelled $\gamma$ so $h_{k}=0$, and we can take $P_{i}$ to be the minimal matching.
Case 5b): Here $b_{\gamma^{\prime}b}=-1$ and $b_{\gamma^{\prime}a}=1$. If $\gamma^{\prime}$ is positive then so is $b$, hence $h_{k}=0$, and we can take $P_{i}$ to be the minimal matching.
Case 6a): Here $b_{\gamma a}=-1$. If $\gamma$ is positive then so is the subsequent diagonal $a$, hence $h_{k}=0$, and we can take $P_{i}$ to be the minimal matching.
Case 6b): If both diagonals $\gamma$ are positive, then so is the diagonal $c$. Thus, recalling the definition of the height function, we have that $h_{k}\geq-1$. Taking $P_{i}$ such that both diagonals $a$ are negative, and all other diagonals are positive achieves $h_{k}=-1$.
Case 7a): There is no diagonal labelled $\gamma$ so $h_{k}=0$, and we can take $P_{i}$ to be the minimal matching.
Case 7b): Here $b_{\gamma^{\prime}b}=-1$ and $b_{\gamma^{\prime}a}=1$. If the diagonal $\gamma^{\prime}$ is positive then so is the diagonal $b$, hence $h_{k}=0$ and we can take $P_{i}$ to be the minimal matching.
Case 8a): Here $b_{\gamma a}=-1$. If the diagonal $\gamma$ is positive then so is the subsequent diagonal $a$, hence $h_{k}=0$ and we can take $P_{i}$ to be the minimal matching.
Case 8b): Here $b_{\gamma^{\prime}a},b_{\gamma^{\prime}c}\leq 0$. Moreover, there is precisely one tile with diagonal $\gamma^{\prime}$, so $h_{k}=-1$. Taking $P_{i}$ such that the diagonal $\gamma^{\prime}$ and the subsequent diagonal $c$ are positive, and all other diagonals negative achieves $h_{k}=-1$.
This completes the proof of Proposition 5.10. ∎
5.6. Proof of equation (5.4) in the Combinatorial Key Lemma
Theorem 5.15.
Equation (5.4) in Theorem 5.3 holds true.
Proof.
Without loss of generality, suppose that the signature of $T$ is non-negative, and identify $T$ with the ideal triangulation $T^{\circ}$ representing it. Consider all the local segments of $\alpha$ with respect to the flipping arc $\gamma:=\tau_{k}$, cf. Subsection 5.3. Some of the arcs lying on the region surrounding these local segments, cf. Figure 8, may be folded sides of self-folded triangles, even the dotted ones. Suppose first that $n>0$ such arcs are folded sides. We let $d_{1},\ldots,d_{m}$ denote these folded sides, allowing repetition. For convenience, we also order $d_{1},\ldots,d_{m}$ in the order that $\alpha$ intersects them (up to cyclic permutation). For each $j=1,\ldots,m$, we let $d_{j}^{\circ}\in T^{\circ}$ denote the loop that together with $d_{j}$ forms the corresponding self-folded triangle of $T^{\circ}$.
The sum
$\sum_{P\in\mathcal{P}(G(T,\alpha))}\overline{y}(P)$
decomposes as the sum of $2^{m}$ collections of perfect matchings – each of these sums corresponding to a choice of orientation on each of the diagonals $d_{1},\ldots,d_{m}$. Some of these sums may be empty; by convention we define an empty sum to be 00. Specifically, if we let $(d_{i_{1}},\ldots,d_{i_{k}})$ denote the collection of good matchings which induce positive orientation on $d_{i_{1}},\ldots,d_{i_{k}}$ and negative orientation on the remaining diagonals, then we have the equality
$\sum_{P\in\mathcal{P}(G(T,\alpha))}\overline{y}(P)=\sum_{(d_{i_{1}},\ldots,d_{i_{k}})}\Big{(}\sum_{P\in(d_{i_{1}},\ldots,d_{i_{k}})}\overline{y}(P)\Big{)}.$
Moreover for any $(d_{i_{1}},\ldots,d_{i_{k}})$ we have
(5.7) $\sum_{P\in(d_{i_{1}},\ldots,d_{i_{k}})}\overline{y}(P)=\Big{(}\prod_{i=1}^{m}\Big{(}\sum_{P\in(d_{i_{1}},\ldots,d_{i_{k}})_{i,i+1}}y(P)\Big{)}\Big{)}\frac{y_{d_{i_{1}}}\ldots y_{d_{i_{k}}}}{y_{d_{i_{1}}^{\circ}}\ldots y_{d_{i_{k}}^{\circ}}},$
where $(d_{i_{1}},\ldots,d_{i_{k}})_{i,i+1}$ denotes the collection of perfect matchings, between the tiles $d_{i}$ and $d_{i+1}$, which induce the orientations on $d_{i}$ and $d_{i+1}$ dictated by $(d_{i_{1}},\ldots,d_{i_{k}})$. By convention we define $(d_{i_{1}},\ldots,d_{i_{k}})_{m,m+1}:=(d_{i_{1}},\ldots,d_{i_{k}})_{m,1}$.
Analogously, if we look at $T^{\prime}$ we get
(5.8) $\sum_{P\in(d_{i_{1}},\ldots,d_{i_{k}})^{\prime}}\overline{y}^{\prime}(P)=\Big{(}\prod_{i=1}^{m}\Big{(}\sum_{P\in(d_{i_{1}},\ldots,d_{i_{k}})^{\prime}_{i,i+1}}y^{\prime}(P)\Big{)}\Big{)}\frac{y_{d_{i_{1}}}^{\prime}\ldots y_{d_{i_{k}}}^{\prime}}{y_{d_{i_{1}}^{\circ}}^{\prime}\ldots y_{d_{i_{k}}^{\circ}}^{\prime}}$
where the collection $(d_{i_{1}},\ldots,d_{i_{k}})^{\prime}$ is defined in the same way as above, but now with respect to $\mathcal{P}(G(T^{\prime},\alpha))$, rather than $\mathcal{P}(G(T,\alpha))$.
By Proposition 5.10 and the explicit computations carried out in Figure 11, the proof of the identity
$(y_{k}+1)^{h_{k}}\cdot\sum_{P\in(d_{i_{1}},\ldots,d_{i_{k}})}\overline{y}(P)=(y_{k}^{\prime}+1)^{h_{k}^{\prime}}\cdot\sum_{P\in(d_{i_{1}},\ldots,d_{i_{k}})^{\prime}}\overline{y}^{\prime}(P)$
is reduced to showing that $\frac{y^{\prime}_{d_{i_{1}}}\ldots y^{\prime}_{d_{i_{k}}}}{y^{\prime}_{d_{i_{1}}^{\circ}}\ldots y^{\prime}_{d_{i_{k}}^{\circ}}}=\frac{y_{d_{i_{1}}}\ldots y_{d_{i_{k}}}}{y_{d_{i_{1}}^{\circ}}\ldots y_{d_{i_{k}}^{\circ}}}.$ This follows from the equality $b_{\gamma\hskip 0.56905ptd_{i}}=b_{\gamma\hskip 0.56905ptd_{i}^{\circ}}$.
Figure 11. “Local” verification of equation (5.4) in the proof of Theorem 5.15.
When none of the boundary edges of the flip regions encountered corresponds to self-folded triangles, the proof of equation (5.4), is easier, because, in that case, the snake subgraphs of the band graphs $G(T,\alpha)$ that arise from the local segment do not interfere with each other at all. The details for that case are left to the reader. ∎
5.7. Local $\mathbf{g}$-vectors
In this section we show that there are six fundamental configurations that dictate how snake g-vectors change under mutation – these are listed below in Figure 12.
Figure 12. Here we list all six combinatorial types of local curves for computing snake g-vectors, together with their corresponding graphs. The red edges arise from the associated minimal matching.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Lemma 5.16.
Let $\alpha$, $T$ and $T^{\prime}$ be as in the statement of Theorem 5.3, and let $\mathcal{L}_{k}$ denote the set of all local curves of $\alpha$ with respect to the flipping/mutating index $k$. Then the following equality holds:
(5.9) $g_{G(T,\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T,\beta)}=g_{G(T^{\prime},\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T^{\prime},\beta)}$
Moreover, the $k^{th}$ component of $g_{G(T,\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T,\beta)}$ is 00.
Proof.
By direct inspection of Figure 12 we see that the graphs
$G(T,\alpha)\setminus\big{(}\displaystyle\coprod_{\beta\in\mathcal{L}_{k}}G(T,\beta)\big{)}\qquad\text{and}\qquad G(T,\alpha)\setminus\big{(}\displaystyle\coprod_{\beta\in\mathcal{L}_{k}}G(T,\beta)\big{)}$
are isomorphic. The validity of Equation (5.9) immediately follows.
Furthermore, let $\gamma\in T$ be the arc corresponding to $k$. Note that $\gamma$ is an edge label of $G(T,\alpha)$ if and only if this edge belongs to $G(T,\beta)$ for some $\beta\in\mathcal{L}_{k}$. Consequently, the $k^{th}$ component of $g_{G(T,\alpha)}-\displaystyle\sum_{\beta\in\mathcal{L}_{k}}g_{G(T,\beta)}$ is 00. ∎
Proposition 5.17.
Let $\alpha$, $T$ and $T^{\prime}$ be as in the statement of Theorem 5.3. Then:
(5.10) $\left[\begin{array}[]{c}-B(T^{\prime})\\ \phantom{-}\mathbf{g}_{G(T^{\prime},\alpha)}\end{array}\right]=\mu_{k}\left(\left[\begin{array}[]{c}-B(T)\\ \phantom{-}\mathbf{g}_{G(T,\alpha)}\end{array}\right]\right).$
Proof.
By Lemma 5.16 it suffices to show equation (5.10) holds for the six configurations listed in Figure 12. The result then follows by direct inspection. Note that a little extra care is needed when treating the scenario that boundary edges of the flip region are labelled by some $\ell_{p}$. To this end, recall that $x_{\ell_{p}}=x_{\beta}x_{\beta^{(p)}}$ for some $\beta\in T$, and let us denote the arc corresponding to $k$ by $\gamma$. The validity of equation (5.10) then follows from the equality $b_{\beta\gamma}=b_{\beta^{(p)}\gamma}$. ∎
Theorem 5.18.
Suppose $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ is a surface with non-empty boundary. Let $\alpha$ be a simple closed curve on $\mathbf{\Sigma}$, and $T$ a tagged triangulation of $\mathbf{\Sigma}$. Then
$\mathbf{g}_{G(T,\alpha)}=\operatorname{Sh}_{T}(\alpha).$
Proof.
By Proposition 4.1, there exists a tagged triangulation $T_{\alpha}$ such that $\operatorname{Sh}_{T_{\alpha}}(\alpha)=\mathbf{g}_{T_{\alpha}}(G(T_{\alpha},\alpha))$. Since $T$ can be obtained from $T_{\alpha}$ by a finite sequence of flips, the theorem follows from Proposition 5.17 and Theorem 3.11. ∎
Remark 5.19.
Theorem 5.18 should be compared to [44, Corollary 6.15-(2)], though the reader is advised to be wary of signs and conventions. Our techniques provide a new, independent proof of Theorem 5.18.
5.8. Proof of equations (5.5) and (5.6) in the Combinatorial Key Lemma
Lemma 5.20.
For each $k\in\{1,\ldots,n\}$ the snake $g$- and $h$-vectors $\mathbf{g}_{G(T,\alpha)}=(g_{\alpha;1},\ldots,g_{\alpha;n})$ and $\mathbf{h}_{G(T,\alpha)}=(h_{\alpha;1},\ldots,h_{\alpha;n})$ of a simple closed curve $\alpha$ satisfy:
(5.11) $h_{\alpha;k}=\min(0,g_{\alpha;k})$
Proof.
From Proposition 5.10 we have:
$h_{\alpha;k}=\displaystyle\sum_{\alpha_{ij}\in\mathcal{L}}h_{\alpha_{ij};k}.$
Directly from the definition of shear coordinates, for any (tagged) triangulation $T$ we have:
$\operatorname{Sh}_{T}(\alpha)=\displaystyle\sum_{\alpha_{ij}\in\mathcal{L}}\operatorname{Sh}_{T}(\alpha_{ij}).$
Moreover, for any $k\in\{1,\ldots,n\}$ the following two statements hold:
  • $g_{\alpha;k}=\operatorname{Sh}_{T}(\alpha;k)$,
  • $\operatorname{Sh}_{T}(\alpha_{ij};k)\geq 0$ for all $\alpha_{ij}\in\mathcal{L}$, or $\operatorname{Sh}_{T}(\alpha_{ij};k)\leq 0$ for all $\alpha_{ij}\in\mathcal{L}$.
Therefore, to prove the lemma it suffices to show
$h_{\alpha_{ij};k}=\min\big{(}0,\operatorname{Sh}_{T}(\alpha_{ij};k)\big{)}$
for every $\beta\in\mathcal{L}$. This follows by direct inspection, and from the explicit computations used in the proof of Proposition 5.10. ∎
Proof of (5.5) and (5.6) from Theorem 5.3..
The validity of these two equations follow directly from Proposition 5.17 and Lemma 5.20. ∎
6. Bangle functions are the generic basis
6.1. The bangle function of a tagged arc belongs to the generic basis
Proposition 6.1.
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with non-empty boundary, and let $\alpha$ be a tagged arc on $\mathbf{\Sigma}$. For every tagged triangulation $T$ of $\mathbf{\Sigma}$, the bangle function $\operatorname{MSW}(G(T,\alpha))$ is equal to the generic value taken by the coefficient-free Caldero-Chapoton function $CC_{A(T)}$ on the irreducible component $\pi_{T}(\alpha)\in\operatorname{DecIrr}^{\tau^{-}}(A(T))$.
Proof.
Since the boundary of $\Sigma$ is not empty, by [15, Proposition 7.10] there exists a sequence of flips that transforms $T$ in a tagged triangulation $T_{\alpha}=f_{k_{m}}\cdots f_{k_{1}}(T)$ containing $\alpha$. The negative simple representation $\mathcal{S}_{\alpha}^{-}(A(T_{\alpha}))$ (see [9, the paragraph preceding Proposition 10.15] or [10, Equation (1.15)]) is the unique point in an irreducible component $Z_{T_{\alpha},\alpha}$ for the Jacobian algebra $A(T_{\alpha})$. This component $Z_{T_{\alpha},\alpha}$ is obviously $\tau^{-}$-rigid, i.e., it is a $\tau^{-}$-reduced component with $E$-invariant zero. By Proposition 3.8 and the invariance of the (injective) $E$-invariant under mutations [10, Theorem 7.1], $Z_{T,\alpha}:=\widetilde{\mu}_{k_{m}}\cdots\widetilde{\mu}_{k_{1}}(Z_{T_{\alpha},\alpha})$ is a $\tau^{-}$-reduced component with $E$-invariant zero for $A(T)$, and the generic value $CC_{A(T)}(Z_{T,\alpha})$ is given by $CC_{A(T)}(\mathcal{M}(T,\alpha))$, where $\mathcal{M}(T,\alpha):=\mu_{k_{m}}\cdots\mu_{k_{1}}(\mathcal{S}_{\alpha}^{-}(A(T_{\alpha})))$.
Now, by [19, Cor. 6.3] and [10, Eqn. (2.14) and Thm. 5.1], $CC_{A(T)}(\mathcal{M}(T,\alpha))$ is the cluster variable that corresponds to $\alpha$ according to [15, Theorem 7.11]. On the other hand, by [43, Theorems 4.9, 4.16 and 4.20], $CC_{A(T)}(\mathcal{M}(T,\alpha))=\operatorname{MSW}(G(T,\alpha))$. ∎
6.2. The bangle function of a closed curve belongs to the generic basis
Proposition 6.2.
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with marked points and with non-empty boundary, and let $\alpha$ be a simple closed curve on $\mathbf{\Sigma}$. For every tagged triangulation $T$ of $\mathbf{\Sigma}$, the bangle function $\operatorname{MSW}(G(T,\alpha))$ is equal to the generic value taken by the coefficient-free Caldero-Chapoton function $CC_{A(T)}$ on the irreducible component $\pi_{T}(\alpha)\in\operatorname{DecIrr}^{\tau^{-}}(A(T))$.
Proof.
Let $T$ be an arbitrary tagged triangulation of $\mathbf{\Sigma}$, and let $T_{\alpha}$, and $M(T_{\alpha},\alpha,\lambda,1)$ be as in Proposition 4.1. By Corollary 4.5, the set
$Z_{T_{\alpha},\alpha}:=\overline{\bigcup_{\lambda\in\mathbb{C}^{*}}\operatorname{GL}_{\mathbf{d}}(\mathbb{C})\cdot M(T_{\alpha},\alpha,\lambda,1)}$
is a generically $\tau^{-}$-reduced indecomposable irreducible component of the representation space $\operatorname{rep}(A(T_{\alpha}),\mathbf{d})$, where $\mathbf{d}:=\underline{\dim}(M(T_{\alpha},\alpha,\lambda,1))$, with
(6.1) $\mathbf{g}_{A(T_{\alpha})}(Z_{T_{\alpha},\alpha})=\mathbf{g}_{G(T_{\alpha},\alpha)}\qquad\text{and}\qquad CC_{A(T_{\alpha})}(Z_{T_{\alpha},\alpha})=\operatorname{MSW}(G(T_{\alpha},\alpha)).$
Thus, $Z_{T_{\alpha},\alpha}=\pi_{T_{\alpha}}(\alpha)$. Moreover, the functions $\mathbf{g}_{A(T_{\alpha})}$ and $CC_{A(T_{\alpha})}$ assume on any point in $\bigcup_{\lambda\in\mathbb{C}^{*}}\operatorname{GL}_{\mathbf{d}}(\mathbb{C})\cdot M(T_{\alpha},\alpha,\lambda,1)$ the values (6.1).
Since the boundary of $\Sigma$ is non-empty, there exists a finite sequence $(T_{0},T_{1},\ldots,T_{m})$ of tagged triangulations, with $T_{0}=T_{\alpha}$ and $T_{m}=T$, such that for each $i=1,\ldots,n$, the triangulation $T_{i}$ is obtained from $T_{i-1}$ by flipping an arc $k_{i}\in T_{i-1}$. By [36, Theorem 8.1] and [9, Proposition 3.7], the Jacobian algebra $A(T):=\mathcal{P}_{\mathbb{C}}(Q^{\prime}(T),S^{\prime}(T))$ is isomorphic to the Jacobian algebra of $\mu_{k_{m}}\cdots\mu_{k_{1}}(Q^{\prime}(T_{\alpha}),S^{\prime}(T_{\alpha}))$. By Proposition 3.8,
$Z_{T,\alpha}:=\widetilde{\mu}_{k_{m}}\cdots\widetilde{\mu}_{k_{1}}(Z_{T_{\alpha},\alpha})$
is a $\tau^{-}$-reduced indecomposable irreducible component of the representation varieties of $\mathcal{P}_{\mathbb{C}}(\mu_{k_{m}}\cdots\mu_{k_{1}}(Q^{\prime}(T_{\alpha}),S^{\prime}(T_{\alpha})))\cong A(T)$. By Theorem 3.12, we have $Z_{T,\alpha}=\pi_{T}(\alpha)$.
Since $CC_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))$ is the generic value $CC_{A(T_{\alpha})}(Z_{T_{\alpha},\alpha})$, Proposition 3.8 implies that $CC_{A(T)}(M(T,\alpha,\lambda,1))$ is the generic value $CC(Z_{T,\alpha})$, where
$M(T,\alpha,\lambda,1):=\mu_{k_{m}}\cdots\mu_{k_{1}}(M(T_{\alpha},\alpha,\lambda,1)).$
By Proposition 4.1 and Corollary 4.5 we have
$\mathbf{g}_{A(T_{\alpha})}(M(T_{\alpha},\alpha,\lambda,1))=\mathbf{g}_{A(T_{\alpha})}(Z_{T_{\alpha},\alpha})=\operatorname{Sh}_{T_{\alpha}}(\alpha)=\mathbf{g}_{G(T_{\alpha},\alpha)},$
so Proposition 3.8 and Theorems 3.12 and 5.18 imply that
$\mathbf{g}_{A(T)}(M(T,\alpha,\lambda,1))=\mathbf{g}_{A(T)}(Z_{T,\alpha})=\operatorname{Sh}_{T}(\alpha)=\mathbf{g}_{G(T,\alpha)}.$
At this point, we have shown that the identity $\mathbf{g}_{A(T^{\prime})}(Z_{T^{\prime},\alpha})=\mathbf{g}_{G(T,\alpha)}$ holds for every tagged triangulation, which allows us to see that Theorem 5.3 and [10, Theorem 5.1 and Lemma 5.2] and the particular equality $F_{M(T_{\alpha},\alpha,\lambda,1)}=F_{G(T_{\alpha},\alpha)}$ that was established in Proposition 4.1, imply that $F_{M(T,\alpha,\lambda,1)}=F_{G(T,\alpha)}$.
We deduce that $CC_{A(T)}(Z_{T,\alpha})=CC_{A(T)}(M(T,\alpha,\lambda,1))=\operatorname{MSW}(G(T,\alpha))$ as elements of the Laurent polynomial ring $\mathbb{Z}[x_{j}^{\pm 1}\ |\ j\in T]$. ∎
6.3. Main result
Recall that for any surface with marked points $\mathbf{\Sigma}$ with a (tagged) triangulation $T$ the Caldero-Chapoton algebra $\mathcal{C}_{A(T)}(\mathbf{\Sigma})$ is spanned by the Caldero-Chapoton functions of all decorated representations of $A(T)$. We have a well-known chain of inclusions $\mathcal{A}(\mathbf{\Sigma})\subset\mathcal{C}_{A(T)}\subset\mathcal{U}(\mathbf{\Sigma})$. See for example [6] for more details. .
Theorem 6.3.
Let $\mathbf{\Sigma}=(\Sigma,\mathbb{M},\mathbb{P})$ be a surface with marked points such that $\partial\Sigma\neq\emptyset$. Then the following holds for each (tagged) triangulation $T$ of $\mathbf{\Sigma}$:
  • (a)
    We have
    • (i)
      $F_{G(T,\alpha)}=F_{\pi_{T}(L)}$ and
    • (ii)
      $\operatorname{MSW}(T,L)=CC_{A(T)}(\pi_{T}(L))$
    for all laminations $L\in\operatorname{Lam}(\mathbf{\Sigma})$, where $\pi_{T}\colon\operatorname{Lam}(\mathbf{\Sigma})\rightarrow\operatorname{DecIrr}^{\tau^{-}}(A(T))$ is the isomorphism of partial KRS-monoids from Theorem 3.12.
  • (b)
    In particular, the set $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)$ of coefficient-free bangle functions is equal to the generic basis $\mathcal{B}_{A(T)}$ of the coefficient-free Caldero-Chapoton algebra $\mathcal{C}_{A(T)}$.
  • (c)
    If $\mathbb{P}=\emptyset$ or $\lvert\mathbb{M}\rvert\geq 2$ the set $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)=\mathcal{B}_{A(T)}$ is a basis of $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$.
Proof.
(a)(i) Note that the case when $L$ consists entirely of tagged arcs is known due to the work of Derksen-Weyman-Zelevinsky [10, Theorem 5.1]. The case when $L$ contains a closed curve follows immediately from Theorem 5.3, Proposition 4.1, and  [10, Lemma 5.2].
(a)(ii) By Theorem 3.12, we know that there is a bijection between the set of single laminates in $\operatorname{Lam}(\mathbf{\Sigma})$ and the set of indecomposable $\tau^{-}$-reduced components of the representation spaces of the Jacobian algebra $A(T)$. Furthermore, we have proved in Propositions 6.1 and 6.2 that for the indecomposable $\tau^{-}$-reduced component $Z_{T,\alpha}=\pi_{T}(\alpha)$ associated to each single laminate $\alpha$, the coefficient-free generic Caldero-Chapoton function $CC_{A(T)}(Z_{T,\alpha})$ is equal to $\operatorname{MSW}(G(T,\alpha))$.
On the other hand, it follows easily from [6, Lemma 4.11] and the definitions, that when $Z=\overline{Z_{1}\oplus Z_{2}}$ for $Z_{1},Z_{2}\in\operatorname{DecIrr}^{\tau^{-}}(A(T)$ (with $E_{A(T)}(Z_{1},Z_{2})=0$) then we have $CC_{A(T)}(Z)=CC_{A(T)}(Z_{1})\cdot CC_{A(T)}(Z_{2})$. Since $\pi_{T}$ is an isomorphism of (tame) partial KRS-monids by Theorem 3.12, our claim follows from these observations and the definition of $\operatorname{MSW}(T,L)$.
(b) Since $\pi_{T}$ is bijective we get by (a) in particular $\mathcal{B}^{\circ}(\mathbf{\Sigma},T)=\mathcal{B}_{A(T)}$. Next, suppose that $\mathbf{\Sigma}$ is not the dreaded torus $\mathbf{\Sigma}_{1}$ (i.e. we exclude the case $g(\Sigma)=1,\mathbb{P}=\emptyset,\mathbb{M}=\{m\}$). Then, since $\partial\Sigma\neq\emptyset$, by [21, Thm. 1.4] the potential $S^{\prime}(T)$ is up to right equivalence the unique non-degenerate potential for $Q^{\prime}(T)$. Thus, in this case $\mathcal{B}_{A(T)}$ is by [22, Cor. 6.14] a basis of the Caldero-Chapoton algebra $\mathcal{C}_{A(T)}$.
In the only remaining case, the dreaded torus, by the main result of [2] we have in particular that $\mathcal{B}^{\circ}(\mathbf{\Sigma}_{1},T)$ is a basis of $\mathcal{A}(\mathbf{\Sigma}_{1})=\mathcal{U}(\mathbf{\Sigma}_{1})$. So our claim follows also in this case since $\mathcal{A}(\mathbf{\Sigma})\subset\mathcal{C}_{A(T)}\subset\mathcal{U}(\mathbf{\Sigma})$.
(c) We have under this hypothesis $\mathcal{A}(\mathbf{\Sigma})=\mathcal{U}(\mathbf{\Sigma})$ either by the main result from [2], or because $\mathcal{A}(\mathbf{\Sigma})$ is locally acyclic by [41, Thm. 10.6 & Thm. 4.1]. So our claim follows from (b) together with the inclusions $\mathcal{A}(\mathbf{\Sigma})\subset\mathcal{C}_{A(T)}\subset\mathcal{U}(\mathbf{\Sigma})$. ∎
7. Concluding remarks and open problems
During the proofs of Theorems 6.1 and 6.2, we have shown that, given $\alpha$, one can find a (tagged) triangulation $T_{\alpha}$ for which one can further find a very concrete (decorated) representation $\mathcal{M}(T_{\alpha},\alpha)$ of the Jacobian algebra $A(T_{\alpha})=\mathcal{P}(Q(T_{\alpha}),S(T_{\alpha}))$ such that
  1. (1)
    $\mathbf{g}_{A(T_{\alpha})}(\mathcal{M}(T_{\alpha},\alpha))=\mathbf{g}_{G(T_{\alpha},\alpha)}$ $F_{M(T_{\alpha},\alpha)}=F_{G(T_{\alpha},\alpha)}.$
  2. (2)
    for any tagged triangulation $T=f_{k_{m}}\cdots f_{k_{1}}(T_{\alpha})$, the decorated representation $\mathcal{M}(T,\alpha):=\mu_{k_{m}}\cdots\mu_{k_{1}}(\mathcal{M}(T_{\alpha},\alpha))$ of the non-degenerate quiver with potential $(Q(T),S(T))$ satisfies $\mathbf{g}_{A(T)}(\mathcal{M}(T,\alpha))=\mathbf{g}_{G(T,\alpha)}$ $F_{M(T,\alpha)}=F_{G(T,\alpha)}.$
Indeed, when $\alpha$ was a tagged arc, we picked $T_{\alpha}$ to be any tagged triangulation containing it, and $\mathcal{M}(T_{\alpha},\alpha)$ to be the negative simple representation of $(Q(T_{\alpha}),S(T_{\alpha}))$ corresponding to $\alpha$; when $\alpha$ was a simple closed curve, we picked $T_{\alpha}$ to be a triangulation such that the Jacobian algebra $\Lambda_{I}$ of the restriction of $(Q(T_{\alpha}),S(T_{\alpha}))$ to the set $I$ of arcs crossed by $\alpha$ is gentle (see Proposition 4.1), and $\mathcal{M}(T_{\alpha},\alpha)$ to be the positive representation of $(Q(T_{\alpha}),S(T_{\alpha}))$ given by any quasi-simple band module $M(T_{\alpha},\alpha,\lambda,1)$ arising from interpreting $\alpha$ as a band on $\Lambda_{I}$.
Now, although $\mathcal{M}(T_{\alpha},\alpha)$ is a well-defined representation, we have not provided an explicit computation of $\mathcal{M}(T,\alpha)$ in general. Of course, when $\mathbb{P}=\varnothing$, the Jacobian algebra $A(T)$ is gentle, and $\mathcal{M}(T,\alpha)$ can be written down explicitly. When $\mathbb{P}\neq\varnothing$ but the signature of $T$ is zero, the Jacobian algebra $A(T)$ is skewed-gentle, and $\mathcal{M}(T,\alpha)$ can be computed explicitly as well, cf. [8, 20, 24]. But when $\mathbb{P}\neq\varnothing$ and $T$ is arbitrary, the decorated representation $\mathcal{M}(T,\alpha)$ still remains to be explicitly computed in general. It should be noticed that the naive candidate for $\mathcal{M}(T,\alpha)$, namely, the obvious string or band representation of the quiver $Q(T)$ induced by $\alpha$, typically fails to be annihilated by the cyclic derivatives of the potential $S(T)$, see e.g. [34, Example 6.2.7]. Explicit computations of $\mathcal{M}(T,\alpha)$ have been carried out by the second named author in [34] in the following situations:
  • when $T$ is a tagged triangulation of positive signature and $\alpha$ is a tagged arc with at most one notch;
  • when $T$ is a tagged triangulation of positive signature and $\alpha$ is a simple closed curve (this is only implicit in [34], but one can check that the results proved therein apply in this situation).
Problem 7.1.
Compute the decorated representation $\mathcal{M}(T,\alpha)$ of $(Q(T),S(T))$ in general. To appreciate the complexity of this task we also direct the reader to  [11] .
On an arguably more important matter, one of the reasons why in this paper we have not considered surfaces with empty boundary whatsoever, is that the proof given in [22] of the linear independence of the set of generic Caldero-Chapoton functions can definitely not be applied for such surfaces.
Problem 7.2.
For a tagged triangulation $T$ of a punctured surface with empty boundary $(\Sigma,\mathbb{P})$, is the set of generic Caldero-Chapoton functions over $\mathcal{P}(Q(T),S(T))$ linearly independent? Is it a basis for the Caldero-Chapoton algebra of $\mathcal{P}(Q(T),S(T))$, or better, for the (upper, coefficient-free) cluster algebra of $\Sigma$? What is its relation to Musiker–Schiffler–Williams’ bangle functions? Here, $S(T)$ is the potential defined in [36].
Acknowledgments
We are grateful to Jan Schröer for many illuminating discussions. The first author acknowledges partial support from PAPIIT grant IN116723 (2023-2025). The work of this paper originated during the third author’s back-to-back visits at IMUNAM courtesy of the first author’s CONACyT-239255 grant and a DGAPA postdoctoral fellowship. He is grateful for the stimulating environment IMUNAM fostered, and for the additional generous support received from the second author’s grants: CONACyT-238754 and a Cátedra Marcos Moshinsky.